### v1

# Generalization and Knowledge Transfer in Abstract Visual Reasoning Models

Abstract
We study generalization and knowledge reuse capabilities of deep neural networks in the domain of abstract visual reasoning (AVR), employing Raven’s Progressive Matrices (RPMs), a recognized benchmark task for assessing AVR abilities. Two knowledge transfer scenarios referring to the I-RAVEN dataset are investigated. Firstly, inspired by generalization assessment capabilities of the PGM dataset and popularity of I-RAVEN, we introduce Attributeless-I-RAVEN, a benchmark with four generalization regimes that allow to test generalization of abstract rules applied to held-out attributes. Secondly, we construct I-RAVEN-Mesh, a dataset that enriches RPMs with a novel component structure comprising line-based patterns, facilitating assessment of progressive knowledge acquisition in transfer learning setting. The developed benchmarks reveal shortcomings of the contemporary deep learning models, which we partly address with Pathways of Normalized Group Convolution (PoNG) model, a novel neural architecture for solving AVR tasks. PoNG excels in both presented challenges, as well as the standard I-RAVEN and PGM setups.
1 Introduction
Generalization, the ability of a model to perform well on unseen data, remains a fundamental challenge in deep learning (DL). While DL methods have demonstrated remarkable achievements in various domains, their generalization capabilities are often questioned, particularly in tasks that demand abstract problem-solving and reasoning skills [6]. One such domain is abstract visual reasoning (AVR) [31, 47, 44, 26] that encompasses tasks requiring (human) fluid intelligence – an aspect of human cognition believed to be crucial for reasoning in never-encountered settings [42, 5]. The most popular AVR tasks are Raven’s Progressive Matrices (RPMs) [38, 39], which constitute a common problem found in human IQ tests. Typical RPMs comprise two components – the context panels arranged in a grid with the bottom-right panel missing and up to 8 answer panels, out of which only one correctly completes the matrix. Solving an RPM instance requires identification of underlying abstract rules applied to certain attributes of the objects composing the instance (see Fig. 1 for an illustrative example).
Design of computational methods capable of tackling RPMs has for decades been an active area of research [8, 12, 10, 24, 22, 45]. Consequently, a number of works considered automatic creation of RPM datasets [30, 48, 29] and a wide suite of predictive models [14, 13] were proposed, with DL methods showing the most promising performance [53, 25]. While this rapid progress led to exceeding the human level in particular problem setups [50, 33], a fundamental challenge of generalization to novel problem settings remains largely unattained.
Initial works designed several RPM datasets [30, 48, 16], however, measuring generalization was not their focus. While some works explored knowledge transfer between related tasks [29, 46], the complexity of the datasets was limited and consequently they didn’t pose a challenge for contemporary DL methods. To measure generalization in modern DL models, the PGM dataset was introduced [2]. PGM defines eight generalization regimes, each specifying the distribution of objects, rules and attributes in train and test splits. For instance, in the Held-out Triples split, a given rule–object–attribute triplet (e.g. Progression on Object’s Size) was assigned only to one of the two splits. In effect, the models were tested on triplet combinations different from training ones, allowing to assess their generalization capabilities. A subsequent work proposed RAVEN [54], another RPM dataset with enriched perceptual complexity of matrices instantiated in seven visual configurations (Center, 2x2Grid, 3x3Grid, Left-Right, Up-Down, Out-InCenter, Out-InGrid). Moreover, the benchmark is characterized by a moderate sample size, i.e. 70K instances, compared to 1.42M RPMs per each of the eight regimes in PGM. Due to this size disparity, subsequent research gravitated towards RAVEN and its revised variants (I-RAVEN [17] and RAVEN-Fair [3]), which didn’t require substantial computational resources to train DL models.
Drawing inspiration from the generalization assessment capabilities of PGM and the broad adoption of RAVEN, this paper proposes a novel suite of generalization challenges stemming from I-RAVEN [17] (a revised variant of RAVEN that removes a bias in RAVEN’s answer panels). The proposed suite of benchmarks allows assessing generalization of AVR models without requiring substantial computational resources.
Firstly, we introduce Attributeless-I-RAVEN, comprising four generalization regimes corresponding to specific held-out attributes ({Position, Type, Size, Color}), resp. Each regime includes four rules ({Constant, Progression, Arithmetic, Distribute Three}) and the training matrices adhere to the Constant rule for the respective attribute, whereas test matrices employ a rule different from Constant for this attribute. In effect, each regime comprises different distributions of training and test data. Note that unlike PGM, in which regimes are oriented over sets of attributes or their range of values, each Attributeless-I-RAVEN regime targets generalization to novel rules concerning a specific attribute, enabling concept-based assessment of generalization capabilities [34, 36].
Secondly, we propose I-RAVEN-Mesh, a variant of I-RAVEN with a new grid-like structure overlaid on the matrices. The dataset enables assessing generalization to incrementally added structures and progressive knowledge acquisition in a transfer learning (TL) setting.
In investigations involving nine contemporary AVR DL models, we observed that the introduced benchmarks present a substantial challenge for the tested methods. This prompted the development of Pathways of Normalized Group Convolution (PoNG), a novel AVR model that excels in both problem setups: generalization to held-out attributes and incremental knowledge acquisition.
Our main contributions can be summarized as follows:
We introduce the Attributeless-I-RAVEN (A-I-RAVEN) dataset that enables measuring generalization across four regimes.
We construct I-RAVEN-Mesh, an extension of I-RAVEN with a new component structure that facilitates assessment of progressive knowledge acquisition in a TL setting.
We evaluate the performance of state-of-the-art AVR models on the introduced benchmarks, uncovering their limitations in terms of generalization to novel problem settings.
We propose a new neural architecture for solving AVR tasks termed PoNG, which excels in addressing both introduced challenges.
2 Related Work
Generalization in AVR.
In recent years, a variety of AVR problems and corresponding datasets have emerged [4, 35, 9, 37, 41, 19, 15, 56] and several attempts have been made to measure generalization in contemporary AVR models based on the introduced benchmarks. In particular, distinct visual configurations were employed in RAVEN to assess how a model trained on one configuration performs on the remaining ones [54, 43, 57]. Although in such a setting the visual aspects of train/test matrices come from different distributions, the underlying rules and attributes remain the same. In contrast, A-I-RAVEN enables studying the generalization of rules applied to held-out attributes, shifting the focus from perception towards reasoning. Besides RPMs, the limits of generalization have been explored in other AVR tasks as well. Visual Analogy Extrapolation Challenge evaluates model’s capacity for extrapolation [49]. However, such specialized datasets might favor models that explicitly embed the notion of extrapolation in their design and aim for being invariant only to specific attributes such as object size or location. Differently, our benchmarks allow verifying the model’s capacity to learn a given concept from the data and generalize it to novel settings. This perspective links our work to the recent literature on concept learning [36, 34]. However, the concept-oriented benchmarks that originate from ARC [6] remain largely unsolved by DL models and pose a significant challenge even for leading multi-modal large language models [32]. In contrast, both benchmarks proposed in this work are attainable by DL models, though further advances in generalization abilities of the models are necessary to consider them solved.
Model architectures.
Preliminary attempts to solve RPMs with DL models involve WReN [2] that reasons over object relations using Relation Network [40], or SRAN [17] that relies on a hierarchical architecture with panel encoders devoted to particular image groups. A common theme enabling generalization in DL models is to explicitly identify RPM objects. To this end, RelBase [43] employs Attend-Infer-Repeat [7], an unsupervised scene decomposition method, STSN [33] utilizes Slot attention [23] to decompose matrix to slots containing particular objects and Temporal Context Normalization (TCN) [49] to normalize latent matrix panel representations in a task-specific context, and MRNet [3] presents a multi-scale architecture. SCL [50] proposes the scattering transformation, CoPINet [55] and CPCNet [52] rely on contrastive architectures, PredRNet [51] learns to minimize the prediction error, and SCAR [28] adapts its computation to the structure of the considered matrix. Despite the high variety of AVR models, experiments on the introduced benchmarks reveal their shortcomings in terms of generalization and knowledge transfer. In this context, we propose PoNG, a new AVR model that excels in the presented tasks by combining parallel architecture, weight sharing, and tactical normalization.
3 Methods
The set of attributes in I-RAVEN is Position, Number, Type, Size, Color and the set of rules is Constant, Progression, Arithmetic, Distribute Three. For an attribute and a dataset split , where train, validation, test, we define the set of rules applicable to in split by . In I-RAVEN all rule–attribute pairs are valid in all splits:
3.1 Attributeless-I-RAVEN
To probe generalization in DL models, we present A-I-RAVEN, a benchmark composed of 4 generalization regimes. Example matrices are illustrated in Fig. 2, with additional samples provided in Appendix A. Each regime, denoted as A/<Attribute> (e.g., A/Type), corresponds to a single held-out attribute. In train and validation splits, all matrices have the Constant rule applied to this attribute. In the test split, the held-out attribute is governed by a rule sampled from . In effect, during training, the model doesn’t see rule–attribute combinations required to solve test matrices. There are no rule-related constraints on the remaining attributes. Since Position and Number attributes are tightly coupled (e.g., it’s impossible to increase object numerosity while keeping their position constant), we allocate a single generalization regime, A/Position, to cover both attributes. In summary, for a held-out attribute , we have:
Attribute | Rule | Description |
Number | Constant | Each image in a given row contains the same number of lines. |
Progression | The count of lines in a given row changes by a constant factor (e.g. ). | |
Arithmetic | The number of lines in the third column is determined based on an arithmetic operation applied to the preceding columns (e.g. ). | |
DT | Three line counts are sampled and spread among images in a given row. | |
Position | Constant | Each image in a given row contains the same position of lines. |
Progression | A panel arrangement is sampled in each row and rotated by degrees in subsequent columns. | |
Arithmetic | The position of lines in the third column is computed based on a set operation (union or difference) applied to the preceding columns. | |
DT | Three line arrangements are sampled and spread among images in a given row. |
3.2 I-RAVEN-Mesh
The other of the proposed benchmarks is designed to probe progressive knowledge acquisition in a TL setting. I-RAVEN-Mesh extends I-RAVEN by introducing a novel visual component overlaid on top of the existing I-RAVEN components (see Fig. 3). Though the dataset can serve as a learning challenge on its own, the main motivation behind its introduction is to employ models pre-trained on I-RAVEN and fine-tune them on I-RAVEN-Mesh with a configurable train sample size, facilitating analysis of their TL performance. The mesh grid comprises from 1 to 12 lines placed in predefined locations. The set of available lines covers the inner and outer edges of a grid (12 lines in total). The mesh component has two attributes: , which govern the count and location of lines, respectively. To each attribute a rule can be applied. Table 1 describes the effect of applying a given rule–attribute pair to the mesh component.
To generate the mesh component of an I-RAVEN-Mesh matrix, we sample one of the two attributes and a corresponding rule that governs its values. As the attributes often depend on each other (e.g., it’s impossible to increase the number of lines while keeping their position constant), we don’t constrain the value of the other attribute. The rule–attribute pairs for the base I-RAVEN components are generated in the same way as in the original dataset. To generate answers to the matrix, we follow the impartial algorithm proposed in I-RAVEN [17]. In addition, each matrix contains at least one incorrect answer that differs from the correct one only in the mesh component, ensuring that the solver has to identify the correct rule governing the mesh component in order to solve the matrix. To facilitate training with an auxiliary loss, in which the model additionally predicts the representation of rules governing the matrix [2], we extend the base set of rule annotations with ones concerning the Mesh component.
3.3 Pathways of Normalized Group Convolution (PoNG)
In initial experiments, we’ve found out that SOTA AVR models struggle in the proposed generalization challenges. Consequently, we introduce PoNG, a novel model that outcompetes baselines across a number of problem settings. The model follows a typical two-stage design. Firstly, it generates an embedding of each image panel. Then, it aggregates representations of matrix panels to predict the index of the correct answer. PoNG is illustrated in Fig. 4 and its details are described in Appendix E.
Let denote an RPM, where is the set of image panels comprising context panels and answer panels , is a greyscale image of height and width , is the one-hot encoded index of the correct answer, is the multi-hot encoded representation of matrix rules of dimensionality using sparse encoding [27]. In each experiment , while is determined by the number of matrix components in the corresponding dataset ( for I-RAVEN-Mesh, otherwise; see Appendix D for details).
Panel encoder.
The first component of the model has the form , where is the input panel embedding of dimensionality . Following RelBase [43], the module comprises blocks of the same architecture. Each block includes parallel pathways that build high-level and low-level features, resp. The first one contains convolutional blocks, each with 2D convolution, ReLU, and Batch Normalization (BN) [18]. The second one contains 2D max pooling followed by 2D convolution. The sum of both pathway results forms the block output. Differently from RelBase, we flatten the height and width dimensions of the resultant embedding, pass it through a linear layer with ReLU, flatten the channel and spatial dimensions, and pass the tensor through a feed-forward residual block with Layer Normalization (LN) [1]. Finally, we concatenate the tensor with a position embedding (a learned -dimensional vector for each cell in the context grid), leading to .
Reasoner.
The second component of the model has the form , where is the panel embedding of ’th answer. For each answer panel, the reasoner produces embedding that describes how well the considered answer fits into the matrix context. Panel embeddings are stacked and processed by a sequence of reasoning blocks interleaved with bottleneck layers for dimensionality reduction. Each reasoning block comprises BN and parallel pathways, outputs of which are added together to form the output of the block. Next, the latent representation is passed through adaptive average pooling, flattened, processed with a linear layer with ReLU, passed through BN and projected with a linear layer to .
Pathways.
The key aspect of the reasoner module are its pathways. Each takes an input tensor of shape , where is the batch size, is the number of channels, and is the feature dimension. In the first reasoning block and corresponds to the number of panel embeddings in the considered group. Pathways are described as follows: P1 – a pointwise 1D convolution layer that mixes panel features at each spatial location; P2 – a sequence of 2 blocks, each comprising 1D convolution, ReLU, and BN, that builds higher level features spanning neighbouring spatial locations; P3 – analogous to P2, but 1D convolution is replaced with a group 1D convolution that splits the tensor into several groups along the channel dimension, applies a 1D convolution with shared weights to each group, and adds together the representations of each group; P4 – analogous to P3, but groups are arranged into pairs concatenated along the channel dimension and processed with a 1D convolution with shared weights. In contrast to [21], the proposed group convolution layers in both P3 and P4 apply TCN [49] to the outputs in each group. In the first layer P3 and P4 split the input tensor into 3 groups, which allows for producing embeddings of each matrix row and each pair of rows, resp. Though we apply the pathways block in the RPM context, we envisage it as a generic module, also applicable to other settings involving a set of vector representations of shape .
Answer prediction.
Eight representations of the context matrix filled-in with the respective answer, , are processed with three prediction heads. The target head employs two linear layers interleaved with ReLU to produce score describing how well the answer aligns with the matrix context. The aggregate rule head computes the sum of inputs and processes it with two linear layers interleaved with ReLU, producing a latent prediction of matrix rules . The target-conditioned rule head processes its input through a linear layer and computes a weighted sum of the resultant embeddings with weights given by the predicted probability distribution over the set of possible answers , where denotes softmax. The model is trained with a joint loss function , where denotes sigmoid, CE cross-entropy, BCE binary cross-entropy, and are balancing coefficients.
4 Experiments
We assess generalization of state-of-the-art models for solving RPMs on A-I-RAVEN, evaluate progressive knowledge acquisition on I-RAVEN-Mesh, and conduct an ablation study to showcase the contribution of the respective modules that constitute PoNG.
Experimental setup.
In all experiments we use the Adam optimizer [20] with , , and a batch size set to . Learning rate is initialized to and reduced -fold (at most 3 times) if no progress is seen in the validation loss in 5 subsequent epochs, and training stops early in the case of 10 epochs without progress. Unless stated otherwise, each model configuration was trained 3 times with a different seed, and we report mean and standard deviation for these runs. In each experiment, we utilize training, validation, and test matrices, following the standard data split protocol taken in prior works [54, 17]. Experiments were run on a worker with a single NVIDIA DGX A100 GPU.
Baselines.
In addition to the simple CNN-LSTM baseline [2], we assess generalization of SOTA AVR models including WReN [2], CoPINet [55], RelBase [43], SCL [50], SRAN [17], CPCNet [52], PredRNet [51], and STSN [33]. For direct comparison, we evaluate all models on I-RAVEN following the above-described experimental setup.
Reproducibility.
To guarantee reproducibility of experiments, we use a fixed set of random seeds and turn off hardware and framework features concerning indeterministic computation wherever possible. Together with the code, we provide the full training script that can be used to run all training jobs. The training job is packaged as a Docker image with fixed dependencies to isolate the configuration of the training environment. The released code allows for generation of all datasets from scratch, eliminating the dependency on file-hosting services required to distribute the data. The code for reproducing all experiments is publicly accessible at: https://github.com/mikomel/raven.
I-RAVEN | Mesh | A/Color | A/Pos. | A/Size | A/Type | |
CPCNet | ||||||
CNN-LSTM | ||||||
CoPINet | ||||||
PredRNet | ||||||
RelBase | ||||||
SCL | ||||||
SRAN | ||||||
STSN | ||||||
WReN | ||||||
PoNG (ours) | 95.9 | 89.3 | 80.3 | 79.3 | 73.5 | 59.4 |
Generalization on Attributeless-I-RAVEN.
In the first set of experiments we evaluate all considered models on four generalization regimes of A-I-RAVEN. The results are presented in Table 2, along with the reference results on I-RAVEN and I-RAVEN-Mesh. PoNG outperforms all selected baselines across all settings. Among baseline models, the best results on A/Color are achieved by RelBase, followed by SCL. In the remaining attributeless regimes, SCL outperforms other baselines with RelBase taking the second place. Interestingly, the top 3 models rely on rather shallow architectures, yet outcompete other methods that rely on a deeper layout, such as SRAN or STSN. This suggests that parameter-efficient AVR models not only excel in solving RPMs but also generalize better.
Generalization regimes of A-I-RAVEN pose a bigger challenge than the base dataset. While PoNG, the best performing model, achieved test accuracy on I-RAVEN, on A-I-RAVEN regimes it scored from (on A/Type) to (on A/Color). Fig. 5 displays the difference in PoNG’s performance on test and validation splits. On I-RAVEN and I-RAVEN-Mesh the difference is negligible, as in these datasets both splits follow the same distribution. However, the difference in attributeless regimes is significant, which indicates the need for further research on generalization. We conclude that A-I-RAVEN provides a suite of challenging regimes of variable complexity, in which even the best performing models are far from solving all test matrices.
Progressive knowledge acquisition on I-RAVEN-Mesh.
In the second set of experiments we employ I-RAVEN-Mesh to examine the TL ability of each model, besides WReN that presented close-to-random performance in all previous tasks. To this end, we consider variants of partial I-RAVEN-Mesh dataset with a fraction of the training set and compare the performance of a model trained from scratch on a partial dataset to that of a model pre-trained on full I-RAVEN and fine-tuned on a part of I-RAVEN-Mesh. Fig. 6 shows that for pre-training CNN-LSTM, CoPINet, SRAN, CPCNet, and RelBase on I-RAVEN leads to gains smaller than p.p., whereas pre-training PoNG, SCL, PredRNet and STSN improved their accuracy by , , and p.p., resp. In addition, TL clearly improved performance of SCL, PredRNet and STSN in all considered settings, in particular for by , and p.p., resp., indicating their capacity for knowledge reuse.
I-RAVEN | Mesh | A/Color | A/Pos. | A/Size | A/Type | |
w/o P3 and P4 | ||||||
w/o TCN | ||||||
union |
Model | Neutral | Inter. | HO-AP | HO-TP | HO-Triples | HO-LT | HO-SC | Extra. |
PredRNet | 97.4 | 70.5 | 63.4 | 67.8 | 23.4 | 27.3 | 13.1 | 19.7 |
SCL | 87.1 | 56.0 | 79.6 | 76.6 | 23.0 | 14.1 | 12.6 | 19.8 |
PoNG (ours) | 98.1 | 75.2 | 92.1 | 97.7 | 46.1 | 16.9 | 12.6 | 19.9 |
Ablation study.
To understand the relevance of particular aspects of PoNG’s design we performed an ablation study considering several simplified model variants. The results are presented in Table 3. The removal of P3 and P4 reduces model performance, especially on A/Type ( p.p.). Similarly, disabling TCN leads to generally worse results, primarily on A/Color ( p.p.) and A/Size ( p.p.). Training without () or () typically reduces model performance, but training with one of these rule-based prediction heads compensates to some degree the lack of the other. However, the removal of both () deteriorates results across all datasets, signifying high relevance of the auxiliary training signal in PoNG’s training.
To confirm the inherent out-of-distribution generalization abilities of PoNG, we evaluated the model on all PGM regimes without performing any hyperparameter optimization (we only changed the batch size to 256 to reduce training time). Table 4 shows that PoNG achieves state-of-the-art results on PGM, particularly on the Held-out Attribute Pairs, Held-out Triple Pairs and Held-out Triples regimes, exceeding the best reference model by , and p.p., resp. We conclude that strong performance of PoNG on A-I-RAVEN and I-RAVEN-Mesh should not be attributed to any specific bias of the model towards these two datasets.
5 Conclusion
We investigate generalization capabilities of DL models in the AVR domain. To accelerate research in this area, we propose two RPM benchmarks. Attributeless-I-RAVEN introduces four generalization regimes of variable complexity that assess model’s capability to solve matrices with rules applied to novel attributes. I-RAVEN-Mesh overlays line-based patterns on top of the RPM, facilitating TL studies. Experiments on nine strong literature AVR models reveal their limitations in terms of generalization. To elevate state-of-the-art, we introduce PoNG, a novel AVR model capitalizing on parallel design, weight sharing, and normalization. PoNG outcompetes all baselines on the presented challenges and achieves significant improvement over SOTA reference models on PGM.
Acknowledgments and Disclosure of Funding
This research was carried out with the support of the Laboratory of Bioinformatics and Computational Genomics and the High Performance Computing Center of the Faculty of Mathematics and Information Science Warsaw University of Technology. Mikołaj Małkiński was funded by the Warsaw University of Technology within the Excellence Initiative: Research University (IDUB) programme. This paper builds on the MSc thesis titled "Transfer learning in abstract visual reasoning domain" by Adam Kowalczyk from the Warsaw University of Technology, Warsaw, Poland.
Appendix A Additional matrix examples
Appendix B Societal impact
Any effective AVR solver could be potentially misused during the online / remote IQ test assessment. Hypothetically, one might use the proposed model to unfairly claim having a higher IQ than it really is. Alternatively, the proposed datasets could be employed to develop such a model. In general, such fraudulent behaviour can lead to unequal evaluation in job recruitment or other competitive situations that relate to IQ scores.
Appendix C Limitations and future work
In this work we study generalization and knowledge transfer in AVR models employing RPM datasets. However, the set of problems in the AVR domain also includes other tasks not covered in the paper [26]. The Machine Number Sense dataset presents visual arithmetic problems [56], VAEC defines an extrapolation challenge [49], while ARC proposes a set of diverse tasks in a few-shot learning setting [6]. Future research in this area may juxtapose the performance of AVR models across a set of benchmarks oriented towards generalization to ensure generalization advances beyond the RPM datasets.
In the paper we claim that the proposed pathways block, a key component of the introduced model, is a generic module also applicable to other tasks that require reasoning over a set of objects (vector embeddings). Nevertheless, the experimental evaluation of PoNG presented in the paper is focused only on RPM benchmarks, including I-RAVEN, I-RAVEN-Mesh, Attributeless-I-RAVEN, and PGM. Assessing model’s performance on other problems constitutes an interesting extension of this work.
Appendix D Dataset details
Rule encoding.
As discussed in Section 3.3, we use sparse encoding [27] to represent the set of matrix rules as a vector , such that for I-RAVEN-Mesh and otherwise. The set of rules in I-RAVEN is {Constant, Progression, Arithmetic, Distribute Three} and the set of attributes is {Position, Number, Type, Size, Color}. It follows that there is unique rule–attribute pairs. In addition, the Left-Right, Up-Down, Out-InCenter, and Out-InGrid configurations in I-RAVEN comprise two components in which rules exist independently, e.g., the Left-Right component contains matrices with separate rules applied to the left and right sides. This gives an upper bound of rule–attribute combinations in each configuration. As discussed in Section 3.2 and presented in Table 1, the Mesh component introduced in I-RAVEN-Mesh comprises two attributes and four rules, leading to a total of rule–attribute combinations per configuration. As an example, in the Up-Down configuration of I-RAVEN-Mesh, there are rule–attribute combinations for the upper component, another for the lower component, and for the Mesh component. The sparse encoding encodes each rule in a matrix as a one-hot vector and applies the OR operation to the set of one-hot vectors, producing a multi-hot representation of matrix rules.
Layer | Hyperparameters |
Conv2D-ReLU-BN2D | |
Conv2D-ReLU-BN2D | |
MaxPool | |
MaxPool | |
Conv2D | |
Sum | |
Conv2D-ReLU-BN2D | |
Conv2D-ReLU-BN2D | |
MaxPool | |
MaxPool | |
Conv2D | |
Sum | |
Flatten (height & width) | |
Linear-ReLU | |
Flatten (depth & spatial) | |
LN | |
Linear-ReLU | |
LN | |
Linear | |
Sum | |
Position embedding |
Layer | Hyperparameters |
Stack | |
BN1D | |
Conv1D | |
Conv1D-ReLU-BN1D | |
G-C-TCN-ReLU-BN1D | |
GP-C-TCN-ReLU-BN1D | |
AvgPool1D | |
BN1D | |
Conv1D | |
Conv1D-ReLU-BN1D | |
Conv1D-ReLU-BN1D | |
G-C-TCN-ReLU-BN1D | |
G-C-TCN-ReLU-BN1D | |
GP-C-TCN-ReLU-BN1D | |
GP-C-TCN-ReLU-BN1D | |
AvgPool1D | |
BN1D | |
Conv1D | |
Conv1D-ReLU-BN1D | |
Conv1D-ReLU-BN1D | |
G-C-TCN-ReLU-BN1D | |
G-C-TCN-ReLU-BN1D | |
GP-C-TCN-ReLU-BN1D | |
GP-C-TCN-ReLU-BN1D | |
Adaptive AvgPool1D | |
Flatten (depth & feature dim) | |
Linear-ReLU-BN1D | |
Linear |
Layer | Hyperparameters |
Target head | |
Linear-ReLU-Linear | |
Linear | |
Aggregate rule head | |
Sum | |
Linear-ReLU | |
Linear | |
Target-conditioned rule head | |
Weighted sum | |
Linear |
Appendix E Model details
Test | Val | Test Val | |
CPCNet | |||
CNN-LSTM | |||
CoPINet | |||
PredRNet | |||
RelBase | |||
SCL | |||
SRAN | |||
STSN | |||
WReN | |||
PoNG (ours) | 95.9 | 95.6 |
Test | Val | Test Val | |
CPCNet | |||
CNN-LSTM | |||
CoPINet | |||
PredRNet | |||
RelBase | |||
SCL | |||
SRAN | |||
STSN | |||
WReN | |||
PoNG (ours) | 89.3 | 89.1 |
Test | Val | Test Val | |
CPCNet | |||
CNN-LSTM | |||
CoPINet | |||
PredRNet | |||
RelBase | |||
SCL | |||
SRAN | |||
STSN | |||
WReN | |||
PoNG (ours) | 80.3 | 96.9 |
Test | Val | Test Val | |
CPCNet | |||
CNN-LSTM | |||
CoPINet | |||
PredRNet | |||
RelBase | |||
SCL | |||
SRAN | |||
STSN | |||
WReN | |||
PoNG (ours) | 79.3 | 98.2 |
Test | Val | Test Val | |
CPCNet | |||
CNN-LSTM | |||
CoPINet | |||
PredRNet | |||
RelBase | |||
SCL | |||
SRAN | |||
STSN | |||
WReN | |||
PoNG (ours) | 73.5 | 97.1 |
Test | Val | Test Val | |
CPCNet | |||
CNN-LSTM | |||
CoPINet | |||
PredRNet | |||
RelBase | |||
SCL | |||
SRAN | |||
STSN | |||
WReN | |||
PoNG (ours) | 59.4 | 96.1 |
Mean | Center | 2x2Grid | 3x3Grid | L-R | U-D | O-IC | O-IG | |
CPCNet | ||||||||
CNN-LSTM | ||||||||
CoPINet | ||||||||
PredRNet | ||||||||
RelBase | ||||||||
SCL | ||||||||
SRAN | ||||||||
STSN | ||||||||
WReN | ||||||||
PoNG (ours) | 95.9 | 99.5 | 97.8 | 91.2 | 98.7 | 98.7 | 98.8 | 86.5 |
Mean | Center | 2x2Grid | 3x3Grid | L-R | U-D | O-IC | O-IG | |
CPCNet | ||||||||
CNN-LSTM | ||||||||
CoPINet | ||||||||
PredRNet | ||||||||
RelBase | 92.5 | |||||||
SCL | ||||||||
SRAN | ||||||||
STSN | ||||||||
WReN | ||||||||
PoNG (ours) | 89.3 | 89.6 | 82.8 | 90.8 | 91.7 | 91.0 | 87.7 |
Mean | Center | 2x2Grid | 3x3Grid | L-R | U-D | O-IC | O-IG | |
CPCNet | ||||||||
CNN-LSTM | ||||||||
CoPINet | ||||||||
PredRNet | ||||||||
RelBase | ||||||||
SCL | ||||||||
SRAN | ||||||||
STSN | ||||||||
WReN | ||||||||
PoNG (ours) | 80.3 | 84.4 | 85.4 | 80.3 | 72.3 | 71.3 | 88.9 | 79.0 |
Mean | Center | 2x2Grid | 3x3Grid | L-R | U-D | O-IC | O-IG | |
CPCNet | ||||||||
CNN-LSTM | ||||||||
CoPINet | ||||||||
PredRNet | 98.9 | |||||||
RelBase | 98.8 | 64.3 | ||||||
SCL | ||||||||
SRAN | ||||||||
STSN | ||||||||
WReN | ||||||||
PoNG (ours) | 79.3 | 99.6 | 54.5 | 44.0 | 98.8 | 98.9 |
Mean | Center | 2x2Grid | 3x3Grid | L-R | U-D | O-IC | O-IG | |
CPCNet | ||||||||
CNN-LSTM | ||||||||
CoPINet | ||||||||
PredRNet | ||||||||
RelBase | ||||||||
SCL | ||||||||
SRAN | ||||||||
STSN | ||||||||
WReN | ||||||||
PoNG (ours) | 73.5 | 84.0 | 81.1 | 75.2 | 67.7 | 65.3 | 78.9 | 62.5 |
Mean | Center | 2x2Grid | 3x3Grid | L-R | U-D | O-IC | O-IG | |
CPCNet | ||||||||
CNN-LSTM | ||||||||
CoPINet | ||||||||
PredRNet | ||||||||
RelBase | ||||||||
SCL | ||||||||
SRAN | ||||||||
STSN | ||||||||
WReN | ||||||||
PoNG (ours) | 59.4 | 58.4 | 69.7 | 65.4 | 56.6 | 56.6 | 53.2 | 55.4 |
Appendix F Extended results
Tables 13 – 13 present the results (mean and standard deviation) of all considered models on test and validation splits and the difference between these two splits. The results support the analysis of dataset difficulty presented in Section 4. The difference in model performance between test and validation splits in I-RAVEN and I-RAVEN-Mesh is negligible. In Attributeless-I-RAVEN regimes, however, the difference is significant, showing limitations of all evaluated models in terms of generalization. The biggest difference concerns the A/Type regime, suggesting that generalization of rules applied to novel shape types constitutes a real challenge for the contemporary models.
Tables 14 – 19 present the results (mean and standard deviation) of all considered models across all matrix configurations. The most challenging configurations in I-RAVEN and I-RAVEN-Mesh are 3x3Grid and Out-InGrid, in which image panels contain more objects than in the remaining configurations. Apparently such setups require stronger reasoning capabilities to correctly identify the rules applied to multiple objects. Also the results on the Left-Right and Up-Down configurations in the A/Color, A/Size, and A/Type regimes are relatively weaker. In these configurations, rules may be applied to both matrix components (left/right and up/down, resp.), which demands stronger reasoning capabilities. This also concerns the Out-InGrid configuration in the A/Position, A/Size, and A/Type regimes, and the Out-InCenter configuration in the A/Type regime. Results in the A/Position regime are close-to-perfect in configurations comprising a single object in each component (Center, Left-Right, Up-Down, and Out-InCenter) and weaker in the remaining configurations (2x2Grid, 3x3Grid and Out-InGrid). This performance drop can be attributed to the fact that Position attribute can only be effectively applied to the 2x2Grid, 3x3Grid and Out-InGrid configurations allowing modification of the object’s position. In the remaining configurations its application does not introduce any changes.
Appendix G Datasheets for datasets
In what follows, we provide the description of the introduced datasets following the Datasheets for Datasets template [11].
Motivation
For what purpose was the dataset created? Was there a specific task in mind? Was there a specific gap that needed to be filled? Please provide a description.
The datasets were created to study generalization and knowledge transfer abilities of AVR models.
Who created this dataset (e.g., which team, research group) and on behalf of which entity (e.g., company, institution, organization)?
The datasets were created by Mikołaj Małkiński, Adam Kowalczyk and Jacek Mańdziuk from the Warsaw University of Technology, Warsaw, Poland.
Who funded the creation of the dataset? If there is an associated grant, please provide the name of the grantor and the grant name and number.
This research was carried out with the support of the Laboratory of Bioinformatics and Computational Genomics and the High Performance Computing Center of the Faculty of Mathematics and Information Science Warsaw University of Technology. Mikołaj Małkiński was funded by the Warsaw University of Technology within the Excellence Initiative: Research University (IDUB) programme.
Any other comments?
None.
Composition
What do the instances that comprise the dataset represent (e.g., documents, photos, people, countries)? Are there multiple types of instances (e.g., movies, users, and ratings; people and interactions between them; nodes and edges)? Please provide a description.
Each dataset instance represents a single Raven’s Progressive Matrix, which is a typical task used in human IQ tests.
How many instances are there in total (of each type, if appropriate)?
Each regime in Attributeless-I-RAVEN as well as the I-RAVEN-Mesh dataset contains instances. The training, validation, and test splits contain , , and matrices, resp. All together there are instances.
Does the dataset contain all possible instances or is it a sample (not necessarily random) of instances from a larger set? If the dataset is a sample, then what is the larger set? Is the sample representative of the larger set (e.g., geographic coverage)? If so, please describe how this representativeness was validated/verified. If it is not representative of the larger set, please describe why not (e.g., to cover a more diverse range of instances, because instances were withheld or unavailable).
The datasets contain a fixed number of instances generated with the data generator. Using a fixed seed ensures reproducibility of the generation process. The data generator allows to configure the number of generated samples.
What data does each instance consist of? “Raw” data (e.g., unprocessed text or images) or features? In either case, please provide a description.
Each RPM instance comprises images that represent the RPM panels, a corresponding index of the correct answer and a representation of rules that govern the matrix. Section 3 of the paper provides additional details. Each instance is packaged as a separate file in the NPZ format, which is a widely-used binary format to store compressed NumPy arrays.
Is there a label or target associated with each instance? If so, please provide a description.
See above.
Is any information missing from individual instances? If so, please provide a description, explaining why this information is missing (e.g., because it was unavailable). This does not include intentionally removed information, but might include, e.g., redacted text.
There’s no missing data.
Are relationships between individual instances made explicit (e.g., users’ movie ratings, social network links)? If so, please describe how these relationships are made explicit.
There are no relationships between individual instances.
Are there recommended data splits (e.g., training, development/validation, testing)? If so, please provide a description of these splits, explaining the rationale behind them.
The datasets are split into training, validation and test splits. Each generated instance contains the split name in its filename, e.g., the RAVEN_1_train.npz file belongs to the train split.
Are there any errors, sources of noise, or redundancies in the dataset? If so, please provide a description.
To the best of our knowledge there are no errors, sources of noise, nor redundancies in the datasets.
Is the dataset self-contained, or does it link to or otherwise rely on external resources (e.g., websites, tweets, other datasets)? If it links to or relies on external resources, a) are there guarantees that they will exist, and remain constant, over time; b) are there official archival versions of the complete dataset (i.e., including the external resources as they existed at the time the dataset was created); c) are there any restrictions (e.g., licenses, fees) associated with any of the external resources that might apply to a future user? Please provide descriptions of all external resources and any restrictions associated with them, as well as links or other access points, as appropriate.
Both datasets are self-contained.
Does the dataset contain data that might be considered confidential (e.g., data that is protected by legal privilege or by doctor-patient confidentiality, data that includes the content of individuals non-public communications)? If so, please provide a description.
No.
Does the dataset contain data that, if viewed directly, might be offensive, insulting, threatening, or might otherwise cause anxiety? If so, please describe why.
No.
Does the dataset relate to people? If not, you may skip the remaining questions in this section.
No.
Does the dataset identify any subpopulations (e.g., by age, gender)? If so, please describe how these subpopulations are identified and provide a description of their respective distributions within the dataset.
N/A.
Is it possible to identify individuals (i.e., one or more natural persons), either directly or indirectly (i.e., in combination with other data) from the dataset? If so, please describe how.
N/A.
Does the dataset contain data that might be considered sensitive in any way (e.g., data that reveals racial or ethnic origins, sexual orientations, religious beliefs, political opinions or union memberships, or locations; financial or health data; biometric or genetic data; forms of government identification, such as social security numbers; criminal history)? If so, please provide a description.
N/A.
Any other comments?
None.
Collection Process
How was the data associated with each instance acquired? Was the data directly observable (e.g., raw text, movie ratings), reported by subjects (e.g., survey responses), or indirectly inferred/derived from other data (e.g., part-of-speech tags, model-based guesses for age or language)? If data was reported by subjects or indirectly inferred/derived from other data, was the data validated/verified? If so, please describe how.
The data was generated with a computer program.
What mechanisms or procedures were used to collect the data (e.g., hardware apparatus or sensor, manual human curation, software program, software API)? How were these mechanisms or procedures validated?
We extended the data generation code used to create I-RAVEN: https://github.com/husheng12345/SRAN. A subset of the dataset was reviewed manually to ensure correctness of the generated matrices.
If the dataset is a sample from a larger set, what was the sampling strategy (e.g., deterministic, probabilistic with specific sampling probabilities)?
The dataset is produced by a generator that creates new RPM instances subject to specified constraints through a pseudo-random process. We use a fixed seed to ensure reproducibility of the generation process.
Who was involved in the data collection process (e.g., students, crowdworkers, contractors) and how were they compensated (e.g., how much were crowdworkers paid)?
The data generator has been written by the authors of this paper without delegating the work to other individuals.
Over what timeframe was the data collected? Does this timeframe match the creation timeframe of the data associated with the instances (e.g., recent crawl of old news articles)? If not, please describe the timeframe in which the data associated with the instances was created.
Development of the datasets lasted from January 2022 to June 2024.
Were any ethical review processes conducted (e.g., by an institutional review board)? If so, please provide a description of these review processes, including the outcomes, as well as a link or other access point to any supporting documentation.
N/A.
Does the dataset relate to people? If not, you may skip the remaining questions in this section.
No.
Did you collect the data from the individuals in question directly, or obtain it via third parties or other sources (e.g., websites)?
N/A.
Were the individuals in question notified about the data collection? If so, please describe (or show with screenshots or other information) how notice was provided, and provide a link or other access point to, or otherwise reproduce, the exact language of the notification itself.
N/A.
Did the individuals in question consent to the collection and use of their data? If so, please describe (or show with screenshots or other information) how consent was requested and provided, and provide a link or other access point to, or otherwise reproduce, the exact language to which the individuals consented.
N/A.
If consent was obtained, were the consenting individuals provided with a mechanism to revoke their consent in the future or for certain uses? If so, please provide a description, as well as a link or other access point to the mechanism (if appropriate).
N/A.
Has an analysis of the potential impact of the dataset and its use on data subjects (e.g., a data protection impact analysis) been conducted? If so, please provide a description of this analysis, including the outcomes, as well as a link or other access point to any supporting documentation.
N/A.
Any other comments?
We bear all responsibility in case of violation of rights.
Preprocessing/cleaning/labeling
Was any preprocessing/cleaning/labeling of the data done (e.g., discretization or bucketing, tokenization, part-of-speech tagging, SIFT feature extraction, removal of instances, processing of missing values)? If so, please provide a description. If not, you may skip the remainder of the questions in this section.
The data created by the generator is ready to be used in a model. No preprocessing, cleaning, or labeling is required.
Was the “raw” data saved in addition to the preprocessed/cleaned/labeled data (e.g., to support unanticipated future uses)? If so, please provide a link or other access point to the “raw” data.
N/A.
Is the software used to preprocess/clean/label the instances available? If so, please provide a link or other access point.
No specific software is required to preprocess, clean, or label the instances. The released code repository contains the code required to reproduce all experiments from the paper, which can be used as a reference implementation for loading the datasets.
Any other comments?
None.
Uses
Has the dataset been used for any tasks already? If so, please provide a description.
The datasets have been used to conduct experiments presented in the paper.
Is there a repository that links to any or all papers or systems that use the dataset? If so, please provide a link or other access point.
N/A.
What (other) tasks could the dataset be used for?
The datasets could be used in a multi-task setting to improve abstract reasoning capabilities of computer vision models.
Is there anything about the composition of the dataset or the way it was collected and preprocessed/cleaned/labeled that might impact future uses? For example, is there anything that a future user might need to know to avoid uses that could result in unfair treatment of individuals or groups (e.g., stereotyping, quality of service issues) or other undesirable harms (e.g., financial harms, legal risks) If so, please provide a description. Is there anything a future user could do to mitigate these undesirable harms?
We do not see any undesirable harms that could apply to future users of the datasets.
Are there tasks for which the dataset should not be used? If so, please provide a description.
The datasets should not be used in human IQ tests, as they were explicitly designed to assess generalization and knowledge transfer abilities of deep learning models.
Any other comments?
None.
Distribution
Will the dataset be distributed to third parties outside of the entity (e.g., company, institution, organization) on behalf of which the dataset was created? If so, please provide a description.
The datasets are publicly available.
How will the dataset will be distributed (e.g., tarball on website, API, GitHub) Does the dataset have a digital object identifier (DOI)?
The datasets are available at the GitHub repository: https://github.com/mikomel/raven. Each dataset has a corresponding DOI: Attributeless-I-RAVEN: https://doi.org/10.57967/hf/2396, I-RAVEN-Mesh: https://doi.org/10.57967/hf/2397.
When will the dataset be distributed?
The datasets are available since June 2024.
Will the dataset be distributed under a copyright or other intellectual property (IP) license, and/or under applicable terms of use (ToU)? If so, please describe this license and/or ToU, and provide a link or other access point to, or otherwise reproduce, any relevant licensing terms or ToU, as well as any fees associated with these restrictions.
The code repository is released under the GPL-3.0 license. This follows the license associated with the generators of the base datasets – RAVEN (https://github.com/WellyZhang/RAVEN) and I-RAVEN (https://github.com/husheng12345/SRAN). The datasets introduced in this paper are released under the CC license.
Have any third parties imposed IP-based or other restrictions on the data associated with the instances? If so, please describe these restrictions, and provide a link or other access point to, or otherwise reproduce, any relevant licensing terms, as well as any fees associated with these restrictions.
No.
Do any export controls or other regulatory restrictions apply to the dataset or to individual instances? If so, please describe these restrictions, and provide a link or other access point to, or otherwise reproduce, any supporting documentation.
No.
Any other comments?
None.
Maintenance
Who will be supporting/hosting/maintaining the dataset?
Mikołaj Małkiński is the maintainer of the code repository.
How can the owner/curator/manager of the dataset be contacted (e.g., email address)?
Email addresses of the corresponding authors are provided on the tile page. In addition, GitHub Issues can be used to conduct public discussions directly in the code repository.
Is there an erratum? If so, please provide a link or other access point.
No.
Will the dataset be updated (e.g., to correct labeling errors, add new instances, delete instances)? If so, please describe how often, by whom, and how updates will be communicated to users (e.g., mailing list, GitHub)?
Future changes will be documented in release notes in the code repository.
If the dataset relates to people, are there applicable limits on the retention of the data associated with the instances (e.g., were individuals in question told that their data would be retained for a fixed period of time and then deleted)? If so, please describe these limits and explain how they will be enforced.
N/A.
Will older versions of the dataset continue to be supported/hosted/maintained? If so, please describe how. If not, please describe how its obsolescence will be communicated to users.
Older versions of the dataset will be available in the history of the code repository.
If others want to extend/augment/build on/contribute to the dataset, is there a mechanism for them to do so? If so, please provide a description. Will these contributions be validated/verified? If so, please describe how. If not, why not? Is there a process for communicating/distributing these contributions to other users? If so, please provide a description.
Contributions are welcome. GitHub Issues of the code repository will be used to communicate between contributors and project maintainers.
Any other comments?
None.

### v2

@@
-Generalization and Knowledge Transfer in Abstract Visual Reasoning Models
+A-I-RAVEN and I-RAVEN-Mesh: Two New Benchmarks for Abstract Visual Reasoning
 Abstract
@@
 Two knowledge transfer scenarios referring to the I-RAVEN dataset are investigated.
-Firstly, inspired by generalization assessment capabilities of the PGM dataset and popularity of I-RAVEN, we introduce Attributeless-I-RAVEN, a benchmark with four generalization regimes that allow to test generalization of abstract rules applied to held-out attributes.
+Firstly, inspired by generalization assessment capabilities of the PGM dataset and popularity of I-RAVEN, we introduce Attributeless-I-RAVEN (A-I-RAVEN), a benchmark with generalization regimes that allow to systematically test generalization of abstract rules applied to held-out attributes at various levels of complexity (primary and extended regimes).
+In contrast to PGM, A-I-RAVEN features compositionality, a variety of figure configurations, and does not require substantial computational resources.
 Secondly, we construct I-RAVEN-Mesh, a dataset that enriches RPMs with a novel component structure comprising line-based patterns, facilitating assessment of progressive knowledge acquisition in transfer learning setting.
-The developed benchmarks reveal shortcomings of the contemporary deep learning models, which we partly address with Pathways of Normalized Group Convolution (PoNG) model, a novel neural architecture for solving AVR tasks.
-PoNG excels in both presented challenges, as well as the standard I-RAVEN and PGM setups.
+We evaluate 13 strong models from the AVR literature on the introduced datasets, revealing their specific shortcomings in generalization and knowledge transfer.
 1 Introduction
 Generalization, the ability of a model to perform well on unseen data, remains a fundamental challenge in deep learning (DL).
-While DL methods have demonstrated remarkable achievements in various domains, their generalization capabilities are often questioned, particularly in tasks that demand abstract problem-solving and reasoning skills [6].
-One such domain is abstract visual reasoning (AVR) [31, 47, 44, 26] that encompasses tasks requiring (human) fluid intelligence – an aspect of human cognition believed to be crucial for reasoning in never-encountered settings [42, 5].
-The most popular AVR tasks are Raven’s Progressive Matrices (RPMs) [38, 39], which constitute a common problem found in human IQ tests.
+While DL methods have demonstrated remarkable achievements in various domains, their generalization capabilities are often questioned, particularly in tasks that demand abstract problem-solving and reasoning skills Chollet (2019).
+One such domain is abstract visual reasoning (AVR) Mitchell (2021); Stabinger et al.
+(2021); van der Maas et al.
+(2021); Małkiński and Mańdziuk (2023) that encompasses tasks requiring (human) fluid intelligence – an aspect of human cognition believed to be crucial for reasoning in never-encountered settings Carpenter et al.
+(1990).
+The most popular AVR tasks are Raven’s Progressive Matrices (RPMs) Raven (1936); Raven and Court (1998), which constitute a common problem found in human IQ tests.
 Typical RPMs comprise two components – the context panels arranged in a grid with the bottom-right panel missing and up to 8 answer panels, out of which only one correctly completes the matrix.
@@
 1 for an illustrative example).
-Design of computational methods capable of tackling RPMs has for decades been an active area of research [8, 12, 10, 24, 22, 45].
-Consequently, a number of works considered automatic creation of RPM datasets [30, 48, 29] and a wide suite of predictive models [14, 13] were proposed, with DL methods showing the most promising performance [53, 25].
-While this rapid progress led to exceeding the human level in particular problem setups [50, 33], a fundamental challenge of generalization to novel problem settings remains largely unattained.
-Initial works designed several RPM datasets [30, 48, 16], however, measuring generalization was not their focus.
-While some works explored knowledge transfer between related tasks [29, 46], the complexity of the datasets was limited and consequently they didn’t pose a challenge for contemporary DL methods.
-To measure generalization in modern DL models, the PGM dataset was introduced [2].
+Design of computational methods capable of tackling RPMs has for decades been an active area of research Evans (1964); Foundalis (2006); Lovett et al.
+(2007); Kunda et al.
+(2010).
+Consequently, a number of works considered automatic creation of RPM datasets Matzen et al.
+(2010); Wang and Su (2015); Mańdziuk and Żychowski (2019) and a wide suite of predictive models Hernández-Orallo et al.
+(2016); Hernández-Orallo (2017) were proposed, with DL methods showing the most promising performance Yang et al.
+(2022); Małkiński and Mańdziuk (2025).
+While this rapid progress led to exceeding the human level in particular problem setups Wu et al.
+(2020); Mondal et al.
+(2023), a fundamental challenge of generalization to novel problem settings remains largely unattained.
+Initial works designed several RPM datasets Matzen et al.
+(2010); Wang and Su (2015); Hoshen and Werman (2017), however, measuring generalization was not their focus.
+While some works explored knowledge transfer between related tasks Mańdziuk and Żychowski (2019); Tomaszewska et al.
+(2022), the complexity of the datasets was limited and consequently they didn’t pose a challenge for contemporary DL methods.
+To measure generalization in modern DL models, the PGM dataset was introduced Barrett et al.
+(2018).
 PGM defines eight generalization regimes, each specifying the distribution of objects, rules and attributes in train and test splits.
@@
 In effect, the models were tested on triplet combinations different from training ones, allowing to assess their generalization capabilities.
-A subsequent work proposed RAVEN [54], another RPM dataset with enriched perceptual complexity of matrices instantiated in seven visual configurations (Center, 2x2Grid, 3x3Grid, Left-Right, Up-Down, Out-InCenter, Out-InGrid).
+A subsequent work proposed RAVEN Zhang et al.
+(2019a), another RPM dataset with enriched perceptual complexity of matrices instantiated in seven visual configurations (Center, 2x2Grid, 3x3Grid, Left-Right, Up-Down, Out-InCenter, Out-InGrid).
 Moreover, the benchmark is characterized by a moderate sample size, i.e.
-70K instances, compared to 1.42M RPMs per each of the eight regimes in PGM.
-Due to this size disparity, subsequent research gravitated towards RAVEN and its revised variants (I-RAVEN [17] and RAVEN-Fair [3]), which didn’t require substantial computational resources to train DL models.
-Drawing inspiration from the generalization assessment capabilities of PGM and the broad adoption of RAVEN, this paper proposes a novel suite of generalization challenges stemming from I-RAVEN [17] (a revised variant of RAVEN that removes a bias in RAVEN’s answer panels).
-The proposed suite of benchmarks allows assessing generalization of AVR models without requiring substantial computational resources.
-Firstly, we introduce Attributeless-I-RAVEN, comprising four generalization regimes corresponding to specific held-out attributes ({Position, Type, Size, Color}), resp.
-Each regime includes four rules ({Constant, Progression, Arithmetic, Distribute Three}) and the training matrices adhere to the Constant rule for the respective attribute, whereas test matrices employ a rule different from Constant for this attribute.
+instances, compared to RPMs per each of the eight regimes in PGM.
+Due to this size disparity, subsequent research gravitated towards RAVEN and its revised variants (I-RAVEN Hu et al.
+(2021) and RAVEN-Fair Benny et al.
+(2021)), which didn’t require substantial computational resources to train DL models.
+Contribution.
+Drawing inspiration from the broad adoption of RAVEN and the generalization assessment capabilities of PGM, this paper proposes a novel suite of generalization challenges stemming from I-RAVEN Hu et al.
+(2021) (a revised variant of RAVEN that removes a bias in RAVEN’s answer panels).
+However, unlike I-RAVEN, the proposed suite of benchmarks allows for a direct assessment of the generalization and knowledge transfer of AVR models.
+Compared to PGM, our datasets feature compositionality and variety of figure configurations, and their processing doesn’t require substantial computational resources.
+Furthermore, they include structural annotations, which are utilized, for example, in recent neuro-symbolic approaches Zhang et al.
+(2021, 2022).
+First, we introduce Attributeless-I-RAVEN (A-I-RAVEN), comprising 10 generalization regimes.
+The primary regimes correspond to specific held-out attributes ({Position, Type, Size, Color}), resp.
+The training matrices in these regimes adhere to the Constant rule for the respective attribute, whereas test matrices employ a rule different from Constant for this attribute (i.e., Progression, Arithmetic, or Distribute Three).
+Moreover, we propose extended regimes: of them feature a held-out attribute pair, while another replace the Constant rule in the training set with each remaining rule.
 In effect, each regime comprises different distributions of training and test data.
-Note that unlike PGM, in which regimes are oriented over sets of attributes or their range of values, each Attributeless-I-RAVEN regime targets generalization to novel rules concerning a specific attribute, enabling concept-based assessment of generalization capabilities [34, 36].
-Secondly, we propose I-RAVEN-Mesh, a variant of I-RAVEN with a new grid-like structure overlaid on the matrices.
+Next, we propose I-RAVEN-Mesh, a variant of I-RAVEN with a new grid-like structure overlaid on the matrices.
 The dataset enables assessing generalization to incrementally added structures and progressive knowledge acquisition in a transfer learning (TL) setting.
-In investigations involving nine contemporary AVR DL models, we observed that the introduced benchmarks present a substantial challenge for the tested methods.
-This prompted the development of Pathways of Normalized Group Convolution (PoNG), a novel AVR model that excels in both problem setups: generalization to held-out attributes and incremental knowledge acquisition.
-Our main contributions can be summarized as follows:
-We introduce the Attributeless-I-RAVEN (A-I-RAVEN) dataset that enables measuring generalization across four regimes.
+Investigations involving contemporary AVR DL models reveal that the introduced benchmarks present a substantial challenge for the tested methods, raising the need for further advancements in this area.
+The key contributions of the paper are summarized below.
+We introduce the A-I-RAVEN dataset that enables measuring generalization across regimes.
 We construct I-RAVEN-Mesh, an extension of I-RAVEN with a new component structure that facilitates assessment of progressive knowledge acquisition in a TL setting.
 We evaluate the performance of state-of-the-art AVR models on the introduced benchmarks, uncovering their limitations in terms of generalization to novel problem settings.
-We propose a new neural architecture for solving AVR tasks termed PoNG, which excels in addressing both introduced challenges.
 2 Related Work
 Generalization in AVR.
-In recent years, a variety of AVR problems and corresponding datasets have emerged [4, 35, 9, 37, 41, 19, 15, 56] and several attempts have been made to measure generalization in contemporary AVR models based on the introduced benchmarks.
-In particular, distinct visual configurations were employed in RAVEN to assess how a model trained on one configuration performs on the remaining ones [54, 43, 57].
+In recent years, a variety of AVR problems and corresponding datasets have emerged Nie et al.
+(2020); Fleuret et al.
+(2011); Qi et al.
+(2021); Shanahan et al.
+(2020); Hill et al.
+(2019); Zhang et al.
+(2020) and several attempts have been made to measure generalization in contemporary AVR models based on the introduced benchmarks.
+In particular, distinct visual configurations were employed in RAVEN to assess how a model trained on one configuration performs on the remaining ones Zhang et al.
+(2019a); Spratley et al.
+(2020); Zhuo and Kankanhalli (2021).
 Although in such a setting the visual aspects of train/test matrices come from different distributions, the underlying rules and attributes remain the same.
@@
 Besides RPMs, the limits of generalization have been explored in other AVR tasks as well.
-Visual Analogy Extrapolation Challenge evaluates model’s capacity for extrapolation [49].
+Visual Analogy Extrapolation Challenge evaluates model’s capacity for extrapolation Webb et al.
+(2020).
 However, such specialized datasets might favor models that explicitly embed the notion of extrapolation in their design and aim for being invariant only to specific attributes such as object size or location.
 Differently, our benchmarks allow verifying the model’s capacity to learn a given concept from the data and generalize it to novel settings.
-This perspective links our work to the recent literature on concept learning [36, 34].
-However, the concept-oriented benchmarks that originate from ARC [6] remain largely unsolved by DL models and pose a significant challenge even for leading multi-modal large language models [32].
+This perspective links our work to the recent literature on concept learning Moskvichev et al.
+(2023).
+However, the concept-oriented benchmarks that originate from ARC Chollet (2019) remain largely unsolved by DL models and pose a significant challenge even for leading multi-modal large language models Mitchell et al.
+(2023).
 In contrast, both benchmarks proposed in this work are attainable by DL models, though further advances in generalization abilities of the models are necessary to consider them solved.
 Model architectures.
-Preliminary attempts to solve RPMs with DL models involve WReN [2] that reasons over object relations using Relation Network [40], or SRAN [17] that relies on a hierarchical architecture with panel encoders devoted to particular image groups.
+Preliminary attempts to solve RPMs with DL models involve WReN Barrett et al.
+(2018) that reasons over object relations using Relation Network Santoro et al.
+(2017), or SRAN Hu et al.
+(2021) that relies on a hierarchical architecture with panel encoders devoted to particular image groups.
 A common theme enabling generalization in DL models is to explicitly identify RPM objects.
-To this end, RelBase [43] employs Attend-Infer-Repeat [7], an unsupervised scene decomposition method, STSN [33] utilizes Slot attention [23] to decompose matrix to slots containing particular objects and Temporal Context Normalization (TCN) [49] to normalize latent matrix panel representations in a task-specific context, and MRNet [3] presents a multi-scale architecture.
-SCL [50] proposes the scattering transformation, CoPINet [55] and CPCNet [52] rely on contrastive architectures, PredRNet [51] learns to minimize the prediction error, and SCAR [28] adapts its computation to the structure of the considered matrix.
+To this end, RelBase Spratley et al.
+(2020) employs Attend-Infer-Repeat, an unsupervised scene decomposition method, STSN Mondal et al.
+(2023) utilizes Slot attention Locatello et al.
+(2020) to decompose matrix to slots containing particular objects and Temporal Context Normalization (TCN) Webb et al.
+(2020) to normalize latent matrix panel representations in a task-specific context, DRNet Zhao et al.
+(2024) relies on a dual-stream design, and MRNet Benny et al.
+(2021) presents a multi-scale architecture.
+SCL Wu et al.
+(2020) proposes the scattering transformation, CoPINet Zhang et al.
+(2019b) and CPCNet Yang et al.
+(2023b) rely on contrastive architectures, PredRNet Yang et al.
+(2023a) learns to minimize the prediction error, ALANS Zhang et al.
+(2021) and PrAE Zhang et al.
+(2022) employ neuro-symbolic architectures, and SCAR Małkiński and Mańdziuk (2024b) adapts its computation to the structure of the considered matrix.
 Despite the high variety of AVR models, experiments on the introduced benchmarks reveal their shortcomings in terms of generalization and knowledge transfer.
-In this context, we propose PoNG, a new AVR model that excels in the presented tasks by combining parallel architecture, weight sharing, and tactical normalization.
-3 Methods
+3 Proposed datasets
 The set of attributes in I-RAVEN is Position, Number, Type, Size, Color and the set of rules is Constant, Progression, Arithmetic, Distribute Three.
-For an attribute and a dataset split , where train, validation, test, we define the set of rules applicable to in split by .
+For attribute and a dataset split , where train, val., test, we define the set of rules applicable to in split by .
 In I-RAVEN all rule–attribute pairs are valid in all splits:
-3.1 Attributeless-I-RAVEN
-To probe generalization in DL models, we present A-I-RAVEN, a benchmark composed of 4 generalization regimes.
+3.1 Attributeless-I-RAVEN (A-I-RAVEN)
+To probe generalization in DL models, we present A-I-RAVEN, a benchmark composed of generalization regimes.
 Example matrices are illustrated in Fig.
 2, with additional samples provided in Appendix A.
-Each regime, denoted as A/<Attribute> (e.g., A/Type), corresponds to a single held-out attribute.
-In train and validation splits, all matrices have the Constant rule applied to this attribute.
-In the test split, the held-out attribute is governed by a rule sampled from .
+Each regime defines a set of held-out attributes , each with a corresponding rule .
+In train and validation splits, held-out attribute is governed by .
+In the test split, is governed by a different rule sampled from .
 In effect, during training, the model doesn’t see rule–attribute combinations required to solve test matrices.
 There are no rule-related constraints on the remaining attributes.
-Since Position and Number attributes are tightly coupled (e.g., it’s impossible to increase object numerosity while keeping their position constant), we allocate a single generalization regime, A/Position, to cover both attributes.
-In summary, for a held-out attribute , we have:
+In summary, we have:
+We define primary regimes with that correspond to individual held-out attributes (), denoted as A/<Attribute> (e.g., A/Type).
+Since Position and Number attributes are tightly coupled (e.g., it’s impossible to increase cardinality of objects while keeping their position constant), we allocate a single generalization regime, A/Position, to cover both attributes.
+In addition, we define extended regimes as supplementary generalization challenges.
+In the first group a pair of attributes is held-out in the training set, i.e.
+.
+Specifically, we introduce new regimes: A/ColorSize, A/ColorType, and A/SizeType, based on the respective attribute pairs.
+In the second group, Constant rule in is replaced with each of the remaining rules, leading to A/Color-Progression, A/Color-Arithmetic, and A/Color-DistributeThree regimes.
+While this modification could be applied to all the described regimes, we focus on the Color attribute due to its broad range of possible values.
 | Attribute | Rule | Description |
@@
 | Arithmetic | The number of lines in the third column is determined based on an arithmetic operation applied to the preceding columns (e.g. ). | |
-DT | Three line counts are sampled and spread among images in a given row. | |
+| Distribute Three | Three line counts are sampled and spread among images in a given row. | |
 | Position | Constant | Each image in a given row contains the same position of lines. |
@@
 | Arithmetic | The position of lines in the third column is computed based on a set operation (union or difference) applied to the preceding columns. | |
-DT | Three line arrangements are sampled and spread among images in a given row. |
+| Distribute Three | Three line arrangements are sampled and spread among images in a given row. |
 3.2 I-RAVEN-Mesh
@@
 Though the dataset can serve as a learning challenge on its own, the main motivation behind its introduction is to employ models pre-trained on I-RAVEN and fine-tune them on I-RAVEN-Mesh with a configurable train sample size, facilitating analysis of their TL performance.
-The mesh grid comprises from 1 to 12 lines placed in predefined locations.
-The set of available lines covers the inner and outer edges of a grid (12 lines in total).
+The mesh grid comprises from to lines placed in predefined locations.
+The set of available lines covers the inner and outer edges of a grid ( lines in total).
 The mesh component has two attributes: , which govern the count and location of lines, respectively.
@@
 The rule–attribute pairs for the base I-RAVEN components are generated in the same way as in the original dataset.
-To generate answers to the matrix, we follow the impartial algorithm proposed in I-RAVEN [17].
+To generate answers to the matrix, we follow the impartial algorithm proposed in I-RAVEN Hu et al.
+(2021).
 In addition, each matrix contains at least one incorrect answer that differs from the correct one only in the mesh component, ensuring that the solver has to identify the correct rule governing the mesh component in order to solve the matrix.
-To facilitate training with an auxiliary loss, in which the model additionally predicts the representation of rules governing the matrix [2], we extend the base set of rule annotations with ones concerning the Mesh component.
-3.3 Pathways of Normalized Group Convolution (PoNG)
-In initial experiments, we’ve found out that SOTA AVR models struggle in the proposed generalization challenges.
-Consequently, we introduce PoNG, a novel model that outcompetes baselines across a number of problem settings.
-The model follows a typical two-stage design.
-Firstly, it generates an embedding of each image panel.
-Then, it aggregates representations of matrix panels to predict the index of the correct answer.
-PoNG is illustrated in Fig.
-4 and its details are described in Appendix E.
-Let denote an RPM, where is the set of image panels comprising context panels and answer panels , is a greyscale image of height and width , is the one-hot encoded index of the correct answer, is the multi-hot encoded representation of matrix rules of dimensionality using sparse encoding [27].
-In each experiment , while is determined by the number of matrix components in the corresponding dataset ( for I-RAVEN-Mesh, otherwise; see Appendix D for details).
-Panel encoder.
-The first component of the model has the form , where is the input panel embedding of dimensionality .
-Following RelBase [43], the module comprises blocks of the same architecture.
-Each block includes parallel pathways that build high-level and low-level features, resp.
-The first one contains convolutional blocks, each with 2D convolution, ReLU, and Batch Normalization (BN) [18].
-The second one contains 2D max pooling followed by 2D convolution.
-The sum of both pathway results forms the block output.
-Differently from RelBase, we flatten the height and width dimensions of the resultant embedding, pass it through a linear layer with ReLU, flatten the channel and spatial dimensions, and pass the tensor through a feed-forward residual block with Layer Normalization (LN) [1].
-Finally, we concatenate the tensor with a position embedding (a learned -dimensional vector for each cell in the context grid), leading to .
-Reasoner.
-The second component of the model has the form , where is the panel embedding of ’th answer.
-For each answer panel, the reasoner produces embedding that describes how well the considered answer fits into the matrix context.
-Panel embeddings are stacked and processed by a sequence of reasoning blocks interleaved with bottleneck layers for dimensionality reduction.
-Each reasoning block comprises BN and parallel pathways, outputs of which are added together to form the output of the block.
-Next, the latent representation is passed through adaptive average pooling, flattened, processed with a linear layer with ReLU, passed through BN and projected with a linear layer to .
-Pathways.
-The key aspect of the reasoner module are its pathways.
-Each takes an input tensor of shape , where is the batch size, is the number of channels, and is the feature dimension.
-In the first reasoning block and corresponds to the number of panel embeddings in the considered group.
-Pathways are described as follows: P1 – a pointwise 1D convolution layer that mixes panel features at each spatial location; P2 – a sequence of 2 blocks, each comprising 1D convolution, ReLU, and BN, that builds higher level features spanning neighbouring spatial locations; P3 – analogous to P2, but 1D convolution is replaced with a group 1D convolution that splits the tensor into several groups along the channel dimension, applies a 1D convolution with shared weights to each group, and adds together the representations of each group; P4 – analogous to P3, but groups are arranged into pairs concatenated along the channel dimension and processed with a 1D convolution with shared weights.
-In contrast to [21], the proposed group convolution layers in both P3 and P4 apply TCN [49] to the outputs in each group.
-In the first layer P3 and P4 split the input tensor into 3 groups, which allows for producing embeddings of each matrix row and each pair of rows, resp.
-Though we apply the pathways block in the RPM context, we envisage it as a generic module, also applicable to other settings involving a set of vector representations of shape .
-Answer prediction.
-Eight representations of the context matrix filled-in with the respective answer, , are processed with three prediction heads.
-The target head employs two linear layers interleaved with ReLU to produce score describing how well the answer aligns with the matrix context.
-The aggregate rule head computes the sum of inputs and processes it with two linear layers interleaved with ReLU, producing a latent prediction of matrix rules .
-The target-conditioned rule head processes its input through a linear layer and computes a weighted sum of the resultant embeddings with weights given by the predicted probability distribution over the set of possible answers , where denotes softmax.
-The model is trained with a joint loss function , where denotes sigmoid, CE cross-entropy, BCE binary cross-entropy, and are balancing coefficients.
+To facilitate training with an auxiliary loss, in which the model additionally predicts the representation of rules governing the matrix Barrett et al.
+(2018), we extend the base set of rule annotations with ones concerning the Mesh component.
+| I-RAVEN† | I-RAVEN (ours) | Mesh | A/Color | A/Position | A/Size | A/Type | |
+| ALANS | |||||||
+| CPCNet | 98.5 | ||||||
+| CNN-LSTM | |||||||
+| CoPINet | |||||||
+| DRNet | 90.9 | 70.0 | 77.5 | ||||
+| MRNet | |||||||
+| PrAE | |||||||
+| PredRNet | |||||||
+| RelBase | 84.9 | ||||||
+| SCL | 65.6 | 49.5 | |||||
+| SRAN | |||||||
+| STSN | |||||||
+| WReN |
+| A/ColorSize | A/ColorType | A/SizeType | A/Color-P | A/Color-A | A/Color-D3 | |
+| ALANS | ||||||
+| CPCNet | ||||||
+| CNN-LSTM | ||||||
+| CoPINet | ||||||
+| DRNet | 66.7 | |||||
+| MRNet | ||||||
+| PrAE | ||||||
+| PredRNet | ||||||
+| RelBase | 65.7 | |||||
+| SCL | 40.8 | 32.0 | 33.5 | 75.6 | ||
+| SRAN | ||||||
+| STSN | ||||||
+| WReN |
 4 Experiments
-We assess generalization of state-of-the-art models for solving RPMs on A-I-RAVEN, evaluate progressive knowledge acquisition on I-RAVEN-Mesh, and conduct an ablation study to showcase the contribution of the respective modules that constitute PoNG.
+We assess generalization of state-of-the-art models for solving RPMs on A-I-RAVEN and evaluate progressive knowledge acquisition on I-RAVEN-Mesh.
 Experimental setup.
-In all experiments we use the Adam optimizer [20] with , , and a batch size set to .
-Learning rate is initialized to and reduced -fold (at most 3 times) if no progress is seen in the validation loss in 5 subsequent epochs, and training stops early in the case of 10 epochs without progress.
-Unless stated otherwise, each model configuration was trained 3 times with a different seed, and we report mean and standard deviation for these runs.
-In each experiment, we utilize training, validation, and test matrices, following the standard data split protocol taken in prior works [54, 17].
+In all experiments we use the Adam optimizer Kingma and Ba (2014) with , , and a batch size set to .
+Learning rate is initialized to and reduced -fold (at most 3 times) if no progress is seen in the validation loss in 5 subsequent epochs, and training stops early in the case of epochs without progress.
+Unless stated otherwise, each model configuration was trained times with a different seed, and we report mean and standard deviation for these runs.
+In each experiment, we utilize training, validation, and test matrices, following the standard data split protocol taken in prior works Zhang et al.
+(2019a); Hu et al.
+(2021).
+All models are trained with the auxiliary loss with sparse encoding Małkiński and Mańdziuk (2024a) and .
 Experiments were run on a worker with a single NVIDIA DGX A100 GPU.
-Baselines.
-In addition to the simple CNN-LSTM baseline [2], we assess generalization of SOTA AVR models including WReN [2], CoPINet [55], RelBase [43], SCL [50], SRAN [17], CPCNet [52], PredRNet [51], and STSN [33].
+Models.
+In addition to the simple CNN-LSTM baseline Barrett et al.
+(2018), we assess generalization of SOTA AVR models including WReN Barrett et al.
+(2018), CoPINet Zhang et al.
+(2019b), RelBase Spratley et al.
+(2020), SCL Wu et al.
+(2020), MRNet Benny et al.
+(2021), ALANS Zhang et al.
+(2021), SRAN Hu et al.
+(2021), PrAE Zhang et al.
+(2022), CPCNet Yang et al.
+(2023b), PredRNet Yang et al.
+(2023a), STSN Mondal et al.
+(2023), and DRNet Zhao et al.
+(2024).
 For direct comparison, we evaluate all models on I-RAVEN following the above-described experimental setup.
@@
 The released code allows for generation of all datasets from scratch, eliminating the dependency on file-hosting services required to distribute the data.
-The code for reproducing all experiments is publicly accessible at: https://github.com/mikomel/raven.
-I-RAVEN | Mesh | A/Color | A/Pos. | A/Size | A/Type | |
-CPCNet | ||||||
-CNN-LSTM | ||||||
-CoPINet | ||||||
-PredRNet | ||||||
-RelBase | ||||||
-SCL | ||||||
-SRAN | ||||||
-STSN | ||||||
-WReN | ||||||
-PoNG (ours) | 95.9 | 89.3 | 80.3 | 79.3 | 73.5 | 59.4 |
-Generalization on Attributeless-I-RAVEN.
-In the first set of experiments we evaluate all considered models on four generalization regimes of A-I-RAVEN.
+The code for reproducing all experiments is publicly accessible at: https://github.com/mikomel/raven
+4.1 Generalization on A-I-RAVEN
+Main regimes.
+In the first set of experiments we evaluate all considered models on primary generalization regimes of A-I-RAVEN.
 The results are presented in Table 2, along with the reference results on I-RAVEN and I-RAVEN-Mesh.
-PoNG outperforms all selected baselines across all settings.
-Among baseline models, the best results on A/Color are achieved by RelBase, followed by SCL.
-In the remaining attributeless regimes, SCL outperforms other baselines with RelBase taking the second place.
-Interestingly, the top 3 models rely on rather shallow architectures, yet outcompete other methods that rely on a deeper layout, such as SRAN or STSN.
-This suggests that parameter-efficient AVR models not only excel in solving RPMs but also generalize better.
-Generalization regimes of A-I-RAVEN pose a bigger challenge than the base dataset.
-While PoNG, the best performing model, achieved test accuracy on I-RAVEN, on A-I-RAVEN regimes it scored from (on A/Type) to (on A/Color).
+The best outcomes on A/Color and A/Position are achieved by DRNet, followed by RelBase and SCL that perform comparably.
+In A/Size and A/Type, SCL outperforms other models, with DRNet and RelBase taking the second and third place, resp.
+Interestingly, the top 3 models present a mix of architectures comprising large models, such as DRNet that includes a Vision Transformer backbone (M params), as well as small models, such as SCL and RelBase that include mainly convolutional and feed-forward layers (M and M params, resp.).
+This suggests that various architectural approaches may be taken to achieve reasonable generalization performance in solving RPMs.
+Extended regimes.
+Table 3 shows the aggregated performance of all considered models on extended A-I-RAVEN regimes.
+Similarly to the main regimes, the best results are achieved by SCL, RelBase, and DRNet.
+Overall, replacing the Constant rule in the training set of the A/Color regime with Progression yields a dataset of slightly lower complexity, as the best model on A/Color-Progression achieved accuracy, a p.p.
+increase compared to the best result on A/Color.
+Conversely, using the Arithmetic and Distribute Three rules increases the difficulty, as measured by the drop of the max accuracy by p.p.
+and p.p., resp.
+Furthermore, using a pair of held-out attributes significantly increases the complexity.
+For instance, in A/ColorType, the most challenging regime, the best result is only .
+We conclude that A-I-RAVEN provides a suite of challenging regimes of variable complexity, in which even the best-performing models are far from solving all test matrices.
+Dataset difficulty.
+Across all A-I-RAVEN regimes, the highest average result was achieved by SCL (), followed by DRNet () and RelBase ().
+While SCL achieved test accuracy on I-RAVEN, on A-I-RAVEN regimes it scored from on A/ColorType to on A/Position.
+Similar differences can be observed for all remaining models, which shows that generalization regimes of A-I-RAVEN pose a bigger challenge than the base dataset.
 Fig.
-5 displays the difference in PoNG’s performance on test and validation splits.
+4 displays the difference in performance of top-3 models on test and validation splits.
 On I-RAVEN and I-RAVEN-Mesh the difference is negligible, as in these datasets both splits follow the same distribution.
-However, the difference in attributeless regimes is significant, which indicates the need for further research on generalization.
-We conclude that A-I-RAVEN provides a suite of challenging regimes of variable complexity, in which even the best performing models are far from solving all test matrices.
-Progressive knowledge acquisition on I-RAVEN-Mesh.
-In the second set of experiments we employ I-RAVEN-Mesh to examine the TL ability of each model, besides WReN that presented close-to-random performance in all previous tasks.
-To this end, we consider variants of partial I-RAVEN-Mesh dataset with a fraction of the training set and compare the performance of a model trained from scratch on a partial dataset to that of a model pre-trained on full I-RAVEN and fine-tuned on a part of I-RAVEN-Mesh.
+However, the difference in attributeless regimes is significant, indicating the need for further research on generalization.
+Tables 4 – 15 in Appendix C present the results of all considered models on test and validation splits and the difference between these two splits for particular datasets/regimes.
+The difference in model performance between test and validation splits in I-RAVEN (Table 4) and I-RAVEN-Mesh (Table 5) is negligible.
+In A-I-RAVEN regimes, however, the difference is significant, showing limitations of all evaluated models in terms of generalization.
+Across primary regimes (Tables 6 – 9), the biggest difference concerns the A/Type regime, suggesting that generalization of rules applied to novel shape types constitutes a real challenge for the contemporary models.
+In all extended regimes concerning held-out attribute pairs (A/ColorSize, A/ColorType, and A/SizeType) the performance difference on test and validation splits is bigger than in the primary regimes (see Tables 10 – 12).
+This drop stems from overall weaker performance on the test split, confirming high difficulty of these regimes.
+Model performance on the next regimes concerning the Color attribute and rules other than Constant (A/Color-Progression, A/Color-Arithmetic, and A/Color-DistributeThree) is better, though further progress in generalization is required to fully close the performance gap between test and validation splits (see Tables 13 – 15).
+Per-configuration results.
+Tables 16 – 27 in Appendix C present the detailed results of all considered models for all matrix configurations.
+The most challenging configurations in I-RAVEN and I-RAVEN-Mesh are 3x3Grid and Out-InGrid, in which image panels contain more objects than in the remaining configurations.
+Apparently, such setups require stronger reasoning capabilities to correctly identify the rules applied to multiple objects.
+Also, the results on the Left-Right and Up-Down configurations are relatively weaker in most regimes.
+In these configurations, rules may be applied to both matrix components (left/right and up/down, resp.), increasing the task complexity.
+This also concerns the Out-InGrid configuration in the A/Size regime, and the Out-InCenter configuration in the A/SizeType regime.
+Results in the A/Position regime are close-to-perfect in configurations comprising a single object in each component (Center, Left-Right, Up-Down, and Out-InCenter) and weaker in the remaining configurations (2x2Grid, 3x3Grid and Out-InGrid).
+This performance drop can be attributed to the fact that Position attribute can only be effectively applied to the 2x2Grid, 3x3Grid and Out-InGrid configurations allowing modification of the object’s position.
+In the remaining configurations its application does not introduce any changes.
+4.2 Progressive knowledge acquisition on I-RAVEN-Mesh
+In the second set of experiments we employ I-RAVEN-Mesh to examine the TL ability of the best performing models.
+To this end, we consider variants of partial I-RAVEN-Mesh dataset with a fraction of the training set and compare the performance of a model trained from scratch on a partial dataset to that of a model pre-trained on full I-RAVEN and fine-tuned on the respective part of I-RAVEN-Mesh.
 Fig.
-6 shows that for pre-training CNN-LSTM, CoPINet, SRAN, CPCNet, and RelBase on I-RAVEN leads to gains smaller than p.p., whereas pre-training PoNG, SCL, PredRNet and STSN improved their accuracy by , , and p.p., resp.
-In addition, TL clearly improved performance of SCL, PredRNet and STSN in all considered settings, in particular for by , and p.p., resp., indicating their capacity for knowledge reuse.
-I-RAVEN | Mesh | A/Color | A/Pos. | A/Size | A/Type | |
-w/o P3 and P4 | ||||||
-w/o TCN | ||||||
-union |
-Model | Neutral | Inter. | HO-AP | HO-TP | HO-Triples | HO-LT | HO-SC | Extra. |
-PredRNet | 97.4 | 70.5 | 63.4 | 67.8 | 23.4 | 27.3 | 13.1 | 19.7 |
-SCL | 87.1 | 56.0 | 79.6 | 76.6 | 23.0 | 14.1 | 12.6 | 19.8 |
-PoNG (ours) | 98.1 | 75.2 | 92.1 | 97.7 | 46.1 | 16.9 | 12.6 | 19.9 |
-Ablation study.
-To understand the relevance of particular aspects of PoNG’s design we performed an ablation study considering several simplified model variants.
-The results are presented in Table 3.
-The removal of P3 and P4 reduces model performance, especially on A/Type ( p.p.).
-Similarly, disabling TCN leads to generally worse results, primarily on A/Color ( p.p.) and A/Size ( p.p.).
-Training without () or () typically reduces model performance, but training with one of these rule-based prediction heads compensates to some degree the lack of the other.
-However, the removal of both () deteriorates results across all datasets, signifying high relevance of the auxiliary training signal in PoNG’s training.
-To confirm the inherent out-of-distribution generalization abilities of PoNG, we evaluated the model on all PGM regimes without performing any hyperparameter optimization (we only changed the batch size to 256 to reduce training time).
-Table 4 shows that PoNG achieves state-of-the-art results on PGM, particularly on the Held-out Attribute Pairs, Held-out Triple Pairs and Held-out Triples regimes, exceeding the best reference model by , and p.p., resp.
-We conclude that strong performance of PoNG on A-I-RAVEN and I-RAVEN-Mesh should not be attributed to any specific bias of the model towards these two datasets.
+5 shows that for pre-training RelBase, MRNet, CPCNet, SRAN, and CoPINet on I-RAVEN leads to gains smaller than p.p., whereas pre-training DRNet, SCL, PredRNet, and STSN improved their accuracy by , , and p.p., resp.
+In addition, TL clearly improved performance of DRNet, SCL, PredRNet, and STSN in all considered settings.
+In particular for by , , , and p.p., resp., indicating the models’ capacity for knowledge reuse.
 5 Conclusion
@@
 To accelerate research in this area, we propose two RPM benchmarks.
-Attributeless-I-RAVEN introduces four generalization regimes of variable complexity that assess model’s capability to solve matrices with rules applied to novel attributes.
+A-I-RAVEN introduces generalization regimes of variable complexity that assess model’s capability to solve matrices with rules applied to novel attributes at various levels of complexity (primary and extended regimes).
+Contrary to the existing PGM dataset, A-I-RAVEN features compositionality, offers a variety of figure configurations, and above all does not require substantial computational resources.
 I-RAVEN-Mesh overlays line-based patterns on top of the RPM, facilitating TL studies.
-Experiments on nine strong literature AVR models reveal their limitations in terms of generalization.
-To elevate state-of-the-art, we introduce PoNG, a novel AVR model capitalizing on parallel design, weight sharing, and normalization.
-PoNG outcompetes all baselines on the presented challenges and achieves significant improvement over SOTA reference models on PGM.
-Acknowledgments and Disclosure of Funding
+Experiments on strong literature AVR models reveal their limitations in terms of generalization.
+We believe that the introduced datasets complement existing RPM benchmarks and will foster progress in the AVR area.
+Limitations and future work.
+In this work we study generalization and knowledge transfer in contemporary AVR models employing RPM datasets.
+While RPMs are by far the most popular AVR tasks, the AVR domain also includes other types of problems not covered in the paper Małkiński and Mańdziuk (2023).
+The Machine Number Sense dataset presents visual arithmetic problems Zhang et al.
+(2020), VAEC defines an extrapolation challenge Webb et al.
+(2020), while ARC proposes a set of diverse tasks in a few-shot learning setting Chollet (2019).
+Similar studies could be performed on problems other than RPMs to test the performance and knowledge transfer abilities of AVR models in other problem settings.
+Acknowledgments
 This research was carried out with the support of the Laboratory of Bioinformatics and Computational Genomics and the High Performance Computing Center of the Faculty of Mathematics and Information Science Warsaw University of Technology.
 Mikołaj Małkiński was funded by the Warsaw University of Technology within the Excellence Initiative: Research University (IDUB) programme.
-This paper builds on the MSc thesis titled "Transfer learning in abstract visual reasoning domain" by Adam Kowalczyk from the Warsaw University of Technology, Warsaw, Poland.
+This paper builds on the MSc thesis titled ”Transfer learning in abstract visual reasoning domain” by Adam Kowalczyk from the Warsaw University of Technology, Warsaw, Poland.
 Appendix A Additional matrix examples
-Appendix B Societal impact
-Any effective AVR solver could be potentially misused during the online / remote IQ test assessment.
-Hypothetically, one might use the proposed model to unfairly claim having a higher IQ than it really is.
-Alternatively, the proposed datasets could be employed to develop such a model.
-In general, such fraudulent behaviour can lead to unequal evaluation in job recruitment or other competitive situations that relate to IQ scores.
-Appendix C Limitations and future work
-In this work we study generalization and knowledge transfer in AVR models employing RPM datasets.
-However, the set of problems in the AVR domain also includes other tasks not covered in the paper [26].
-The Machine Number Sense dataset presents visual arithmetic problems [56], VAEC defines an extrapolation challenge [49], while ARC proposes a set of diverse tasks in a few-shot learning setting [6].
-Future research in this area may juxtapose the performance of AVR models across a set of benchmarks oriented towards generalization to ensure generalization advances beyond the RPM datasets.
-In the paper we claim that the proposed pathways block, a key component of the introduced model, is a generic module also applicable to other tasks that require reasoning over a set of objects (vector embeddings).
-Nevertheless, the experimental evaluation of PoNG presented in the paper is focused only on RPM benchmarks, including I-RAVEN, I-RAVEN-Mesh, Attributeless-I-RAVEN, and PGM.
-Assessing model’s performance on other problems constitutes an interesting extension of this work.
-Appendix D Dataset details
-Rule encoding.
-As discussed in Section 3.3, we use sparse encoding [27] to represent the set of matrix rules as a vector , such that for I-RAVEN-Mesh and otherwise.
+Figure 6 presents matrix examples from A/Type and A/Size, the primary regimes of A-I-RAVEN.
+Figures 7, 8 and 9 depict matrix examples from the extended regimes of A-I-RAVEN: A/ColorSize and A/ColorType (Fig.
+7), A/SizeType and A/Color-Progression (Fig.
+8), and A/Color-Arithmetic and A/Color-DistributeThree (Fig.
+9).
+Figure 10 presents matrix examples from I-RAVEN-Mesh concerning the Number attribute.
+Appendix B Auxiliary training
+As discussed in Section 4, we use sparse encoding Małkiński and Mańdziuk [2024a] to represent the set of matrix rules as a vector , such that for I-RAVEN-Mesh and otherwise.
 The set of rules in I-RAVEN is {Constant, Progression, Arithmetic, Distribute Three} and the set of attributes is {Position, Number, Type, Size, Color}.
@@
 The sparse encoding encodes each rule in a matrix as a one-hot vector and applies the OR operation to the set of one-hot vectors, producing a multi-hot representation of matrix rules.
-Layer | Hyperparameters |
-Conv2D-ReLU-BN2D | |
-Conv2D-ReLU-BN2D | |
-MaxPool | |
-MaxPool | |
-Conv2D | |
-Sum | |
-Conv2D-ReLU-BN2D | |
-Conv2D-ReLU-BN2D | |
-MaxPool | |
-MaxPool | |
-Conv2D | |
-Sum | |
-Flatten (height & width) | |
-Linear-ReLU | |
-Flatten (depth & spatial) | |
-LN | |
-Linear-ReLU | |
-LN | |
-Linear | |
-Sum | |
-Position embedding |
-Layer | Hyperparameters |
-Stack | |
-BN1D | |
-Conv1D | |
-Conv1D-ReLU-BN1D | |
-G-C-TCN-ReLU-BN1D | |
-GP-C-TCN-ReLU-BN1D | |
-AvgPool1D | |
-BN1D | |
-Conv1D | |
-Conv1D-ReLU-BN1D | |
-Conv1D-ReLU-BN1D | |
-G-C-TCN-ReLU-BN1D | |
-G-C-TCN-ReLU-BN1D | |
-GP-C-TCN-ReLU-BN1D | |
-GP-C-TCN-ReLU-BN1D | |
-AvgPool1D | |
-BN1D | |
-Conv1D | |
-Conv1D-ReLU-BN1D | |
-Conv1D-ReLU-BN1D | |
-G-C-TCN-ReLU-BN1D | |
-G-C-TCN-ReLU-BN1D | |
-GP-C-TCN-ReLU-BN1D | |
-GP-C-TCN-ReLU-BN1D | |
-Adaptive AvgPool1D | |
-Flatten (depth & feature dim) | |
-Linear-ReLU-BN1D | |
-Linear |
-Layer | Hyperparameters |
-Target head | |
-Linear-ReLU-Linear | |
-Linear | |
-Aggregate rule head | |
-Sum | |
-Linear-ReLU | |
-Linear | |
-Target-conditioned rule head | |
-Weighted sum | |
-Linear |
-Appendix E Model details
+Appendix C Extended results
+Tables 4 – 15 present the results (mean and standard deviation) of all considered models on test and validation splits and the difference between these two splits for particular datasets/regimes.
+Tables 16 – 27 present the results (mean and standard deviation) of all considered models in detail for all matrix configurations.
 | Test | Val | Test Val | |
+| ALANS | |||
 | CPCNet | |||
@@
 | CoPINet | |||
+| DRNet | 90.9 | 90.8 | |
+| MRNet | |||
+| PrAE | |||
 | PredRNet | |||
@@
 | STSN | |||
-WReN | |||
-PoNG (ours) | 95.9 | 95.6 |
+| WReN |
 | Test | Val | Test Val | |
+| ALANS | |||
 | CPCNet | |||
@@
 | CoPINet | |||
+| DRNet | |||
+| MRNet | |||
+| PrAE | |||
+| PredRNet | |||
+| RelBase | 84.9 | 85.0 | |
+| SCL | |||
+| SRAN | |||
+| STSN | |||
+| WReN |
+| Test | Val | Test Val | |
+| ALANS | |||
+| CPCNet | |||
+| CNN-LSTM | |||
+| CoPINet | |||
+| DRNet | 70.0 | 95.4 | |
+| MRNet | |||
+| PrAE | |||
 | PredRNet | |||
@@
 | STSN | |||
-WReN | |||
-PoNG (ours) | 89.3 | 89.1 |
+| WReN |
 | Test | Val | Test Val | |
+| ALANS | |||
 | CPCNet | |||
@@
 | CoPINet | |||
+| DRNet | 77.5 | 97.8 | |
+| MRNet | |||
+| PrAE | |||
 | PredRNet | |||
@@
 | STSN | |||
-WReN | |||
-PoNG (ours) | 80.3 | 96.9 |
+| WReN |
 | Test | Val | Test Val | |
+| ALANS | |||
 | CPCNet | |||
@@
 | CoPINet | |||
+| DRNet | |||
+| MRNet | |||
+| PrAE | |||
+| PredRNet | |||
+| RelBase | |||
+| SCL | 65.6 | 94.0 | |
+| SRAN | |||
+| STSN | |||
+| WReN |
+| Test | Val | Test Val | |
+| ALANS | |||
+| CPCNet | |||
+| CNN-LSTM | |||
+| CoPINet | |||
+| DRNet | |||
+| MRNet | |||
+| PrAE | |||
+| PredRNet | |||
+| RelBase | |||
+| SCL | 49.5 | 95.9 | |
+| SRAN | |||
+| STSN | |||
+| WReN |
+| Test | Val | Test Val | |
+| ALANS | |||
+| CPCNet | |||
+| CNN-LSTM | |||
+| CoPINet | |||
+| DRNet | 95.9 | ||
+| MRNet | |||
+| PrAE | |||
+| PredRNet | |||
+| RelBase | |||
+| SCL | 40.8 | ||
+| SRAN | |||
+| STSN | |||
+| WReN |
+| Test | Val | Test Val | |
+| ALANS | |||
+| CPCNet | |||
+| CNN-LSTM | |||
+| CoPINet | |||
+| DRNet | |||
+| MRNet | |||
+| PrAE | |||
+| PredRNet | |||
+| RelBase | |||
+| SCL | 32.0 | 96.4 | |
+| SRAN | |||
+| STSN | |||
+| WReN |
+| Test | Val | Test Val | |
+| ALANS | |||
+| CPCNet | |||
+| CNN-LSTM | |||
+| CoPINet | |||
+| DRNet | |||
+| MRNet | 97.7 | ||
+| PrAE | |||
+| PredRNet | |||
+| RelBase | |||
+| SCL | 33.5 | ||
+| SRAN | |||
+| STSN | |||
+| WReN |
+| Test | Val | Test Val | |
+| ALANS | |||
+| CPCNet | |||
+| CNN-LSTM | |||
+| CoPINet | |||
+| DRNet | 96.1 | ||
+| MRNet | |||
+| PrAE | |||
+| PredRNet | |||
+| RelBase | |||
+| SCL | 75.6 | ||
+| SRAN | |||
+| STSN | |||
+| WReN |
+| Test | Val | Test Val | |
+| ALANS | |||
+| CPCNet | |||
+| CNN-LSTM | |||
+| CoPINet | |||
+| DRNet | 66.7 | 93.5 | |
+| MRNet | |||
+| PrAE | |||
 | PredRNet | |||
@@
 | STSN | |||
-WReN | |||
-PoNG (ours) | 79.3 | 98.2 |
+| WReN |
 | Test | Val | Test Val | |
+| ALANS | |||
 | CPCNet | |||
@@
 | CoPINet | |||
+| DRNet | 95.1 | ||
+| MRNet | |||
+| PrAE | |||
 | PredRNet | |||
-RelBase | |||
+| RelBase | 65.7 | ||
 | SCL | |||
@@
 | STSN | |||
-WReN | |||
-PoNG (ours) | 73.5 | 97.1 |
-Test | Val | Test Val | |
-CPCNet | |||
-CNN-LSTM | |||
-CoPINet | |||
-PredRNet | |||
-RelBase | |||
-SCL | |||
-SRAN | |||
-STSN | |||
-WReN | |||
-PoNG (ours) | 59.4 | 96.1 |
+| WReN |
 | Mean | Center | 2x2Grid | 3x3Grid | L-R | U-D | O-IC | O-IG | |
+| ALANS | ||||||||
 | CPCNet | ||||||||
@@
 | CoPINet | ||||||||
-PredRNet | ||||||||
+| DRNet | 90.9 | 83.6 | 98.2 | 98.1 | 70.3 | |||
+| MRNet | 99.8 | |||||||
+| PrAE | ||||||||
+| PredRNet | 93.4 | 98.5 | ||||||
 | RelBase | ||||||||
@@
 | STSN | ||||||||
-WReN | ||||||||
-PoNG (ours) | 95.9 | 99.5 | 97.8 | 91.2 | 98.7 | 98.7 | 98.8 | 86.5 |
+| WReN |
 | Mean | Center | 2x2Grid | 3x3Grid | L-R | U-D | O-IC | O-IG | |
+| ALANS | ||||||||
 | CPCNet | ||||||||
@@
 | CoPINet | ||||||||
+| DRNet | 93.7 | 76.0 | ||||||
+| MRNet | 90.1 | 91.2 | ||||||
+| PrAE | ||||||||
 | PredRNet | ||||||||
-RelBase | 92.5 | |||||||
+| RelBase | 84.9 | 81.5 | 75.5 | 90.3 | ||||
 | SCL | ||||||||
@@
 | STSN | ||||||||
-WReN | ||||||||
-PoNG (ours) | 89.3 | 89.6 | 82.8 | 90.8 | 91.7 | 91.0 | 87.7 |
+| WReN |
 | Mean | Center | 2x2Grid | 3x3Grid | L-R | U-D | O-IC | O-IG | |
+| ALANS | ||||||||
 | CPCNet | ||||||||
@@
 | CoPINet | ||||||||
+| DRNet | 70.0 | 63.9 | 68.1 | 70.1 | 82.7 | |||
+| MRNet | ||||||||
+| PrAE | ||||||||
 | PredRNet | ||||||||
+| RelBase | 66.8 | |||||||
+| SCL | 71.3 | 73.5 | ||||||
+| SRAN | ||||||||
+| STSN | ||||||||
+| WReN |
+| Mean | Center | 2x2Grid | 3x3Grid | L-R | U-D | O-IC | O-IG | |
+| ALANS | ||||||||
+| CPCNet | ||||||||
+| CNN-LSTM | ||||||||
+| CoPINet | ||||||||
+| DRNet | 77.5 | 99.4 | 99.2 | 65.4 | ||||
+| MRNet | ||||||||
+| PrAE | 56.5 | 50.2 | ||||||
+| PredRNet | 98.7 | 98.9 | ||||||
 | RelBase | ||||||||
@@
 | STSN | ||||||||
-WReN | ||||||||
-PoNG (ours) | 80.3 | 84.4 | 85.4 | 80.3 | 72.3 | 71.3 | 88.9 | 79.0 |
+| WReN |
 | Mean | Center | 2x2Grid | 3x3Grid | L-R | U-D | O-IC | O-IG | |
+| ALANS | ||||||||
 | CPCNet | ||||||||
@@
 | CoPINet | ||||||||
-PredRNet | 98.9 | |||||||
-RelBase | 98.8 | 64.3 | ||||||
-SCL | ||||||||
+| DRNet | ||||||||
+| MRNet | ||||||||
+| PrAE | ||||||||
+| PredRNet | ||||||||
+| RelBase | ||||||||
+| SCL | 65.6 | 66.2 | 72.1 | 67.0 | 61.7 | 62.0 | 68.6 | 61.5 |
 | SRAN | ||||||||
 | STSN | ||||||||
-WReN | ||||||||
-PoNG (ours) | 79.3 | 99.6 | 54.5 | 44.0 | 98.8 | 98.9 |
+| WReN |
 | Mean | Center | 2x2Grid | 3x3Grid | L-R | U-D | O-IC | O-IG | |
+| ALANS | ||||||||
 | CPCNet | ||||||||
@@
 | CoPINet | ||||||||
+| DRNet | 47.3 | |||||||
+| MRNet | ||||||||
+| PrAE | ||||||||
 | PredRNet | ||||||||
 | RelBase | ||||||||
-SCL | ||||||||
+| SCL | 49.5 | 58.8 | 57.6 | 47.8 | 47.2 | 40.9 | ||
 | SRAN | ||||||||
-STSN | ||||||||
-WReN | ||||||||
-PoNG (ours) | 73.5 | 84.0 | 81.1 | 75.2 | 67.7 | 65.3 | 78.9 | 62.5 |
+| STSN | 48.5 | |||||||
+| WReN |
 | Mean | Center | 2x2Grid | 3x3Grid | L-R | U-D | O-IC | O-IG | |
+| ALANS | ||||||||
 | CPCNet | ||||||||
@@
 | CoPINet | ||||||||
+| DRNet | 33.0 | |||||||
+| MRNet | ||||||||
+| PrAE | ||||||||
+| PredRNet | ||||||||
+| RelBase | 37.1 | |||||||
+| SCL | 40.8 | 40.8 | 49.2 | 45.7 | 43.4 | 46.0 | ||
+| SRAN | ||||||||
+| STSN | ||||||||
+| WReN |
+| Mean | Center | 2x2Grid | 3x3Grid | L-R | U-D | O-IC | O-IG | |
+| ALANS | ||||||||
+| CPCNet | ||||||||
+| CNN-LSTM | ||||||||
+| CoPINet | ||||||||
+| DRNet | 38.0 | |||||||
+| MRNet | ||||||||
+| PrAE | 30.3 | 28.1 | ||||||
+| PredRNet | ||||||||
+| RelBase | 27.5 | |||||||
+| SCL | 32.0 | 25.9 | 44.4 | 40.1 | ||||
+| SRAN | ||||||||
+| STSN | ||||||||
+| WReN |
+| Mean | Center | 2x2Grid | 3x3Grid | L-R | U-D | O-IC | O-IG | |
+| ALANS | ||||||||
+| CPCNet | ||||||||
+| CNN-LSTM | ||||||||
+| CoPINet | ||||||||
+| DRNet | 29.4 | |||||||
+| MRNet | ||||||||
+| PrAE | ||||||||
 | PredRNet | ||||||||
 | RelBase | ||||||||
-SCL | ||||||||
+| SCL | 33.5 | 41.3 | 44.0 | 39.7 | 30.7 | 29.3 | 23.2 | |
 | SRAN | ||||||||
 | STSN | ||||||||
-WReN | ||||||||
-PoNG (ours) | 59.4 | 58.4 | 69.7 | 65.4 | 56.6 | 56.6 | 53.2 | 55.4 |
-Appendix F Extended results
-Tables 13 – 13 present the results (mean and standard deviation) of all considered models on test and validation splits and the difference between these two splits.
-The results support the analysis of dataset difficulty presented in Section 4.
-The difference in model performance between test and validation splits in I-RAVEN and I-RAVEN-Mesh is negligible.
-In Attributeless-I-RAVEN regimes, however, the difference is significant, showing limitations of all evaluated models in terms of generalization.
-The biggest difference concerns the A/Type regime, suggesting that generalization of rules applied to novel shape types constitutes a real challenge for the contemporary models.
-Tables 14 – 19 present the results (mean and standard deviation) of all considered models across all matrix configurations.
-The most challenging configurations in I-RAVEN and I-RAVEN-Mesh are 3x3Grid and Out-InGrid, in which image panels contain more objects than in the remaining configurations.
-Apparently such setups require stronger reasoning capabilities to correctly identify the rules applied to multiple objects.
-Also the results on the Left-Right and Up-Down configurations in the A/Color, A/Size, and A/Type regimes are relatively weaker.
-In these configurations, rules may be applied to both matrix components (left/right and up/down, resp.), which demands stronger reasoning capabilities.
-This also concerns the Out-InGrid configuration in the A/Position, A/Size, and A/Type regimes, and the Out-InCenter configuration in the A/Type regime.
-Results in the A/Position regime are close-to-perfect in configurations comprising a single object in each component (Center, Left-Right, Up-Down, and Out-InCenter) and weaker in the remaining configurations (2x2Grid, 3x3Grid and Out-InGrid).
-This performance drop can be attributed to the fact that Position attribute can only be effectively applied to the 2x2Grid, 3x3Grid and Out-InGrid configurations allowing modification of the object’s position.
-In the remaining configurations its application does not introduce any changes.
-Appendix G Datasheets for datasets
-In what follows, we provide the description of the introduced datasets following the Datasheets for Datasets template [11].
+| WReN |
+| Mean | Center | 2x2Grid | 3x3Grid | L-R | U-D | O-IC | O-IG | |
+| ALANS | ||||||||
+| CPCNet | ||||||||
+| CNN-LSTM | ||||||||
+| CoPINet | ||||||||
+| DRNet | 85.4 | 77.6 | ||||||
+| MRNet | ||||||||
+| PrAE | 75.0 | 83.0 | 82.9 | |||||
+| PredRNet | ||||||||
+| RelBase | ||||||||
+| SCL | 75.6 | 84.0 | 66.4 | |||||
+| SRAN | ||||||||
+| STSN | ||||||||
+| WReN |
+| Mean | Center | 2x2Grid | 3x3Grid | L-R | U-D | O-IC | O-IG | |
+| ALANS | ||||||||
+| CPCNet | ||||||||
+| CNN-LSTM | ||||||||
+| CoPINet | ||||||||
+| DRNet | 66.7 | 62.5 | 63.3 | 63.2 | 77.5 | |||
+| MRNet | ||||||||
+| PrAE | ||||||||
+| PredRNet | ||||||||
+| RelBase | 69.4 | 72.0 | ||||||
+| SCL | 62.8 | |||||||
+| SRAN | ||||||||
+| STSN | ||||||||
+| WReN |
+| Mean | Center | 2x2Grid | 3x3Grid | L-R | U-D | O-IC | O-IG | |
+| ALANS | ||||||||
+| CPCNet | ||||||||
+| CNN-LSTM | ||||||||
+| CoPINet | ||||||||
+| DRNet | 58.2 | |||||||
+| MRNet | ||||||||
+| PrAE | 69.6 | 69.5 | ||||||
+| PredRNet | ||||||||
+| RelBase | 65.7 | 64.9 | 75.6 | |||||
+| SCL | 70.8 | 71.8 | ||||||
+| SRAN | ||||||||
+| STSN | ||||||||
+| WReN |
+Appendix D Datasheets for datasets
+In what follows, we provide the description of the introduced datasets following the Datasheets for Datasets template proposed by Gebru et al. Gebru et al. [2021].
 Motivation
@@
 How many instances are there in total (of each type, if appropriate)?
-Each regime in Attributeless-I-RAVEN as well as the I-RAVEN-Mesh dataset contains instances. The training, validation, and test splits contain , , and matrices, resp. All together there are instances.
+Each regime in A-I-RAVEN as well as the I-RAVEN-Mesh dataset contains instances. The training, validation, and test splits contain , , and matrices, resp. All together there are () instances.
 Does the dataset contain all possible instances or is it a sample (not necessarily random) of instances from a larger set?
@@
 If not, please describe the timeframe in which the data associated with the instances was created.
-Development of the datasets lasted from January 2022 to June 2024.
+Development of the datasets lasted from January 2022 to September 2024.
 Were any ethical review processes conducted (e.g., by an institutional review board)?